### v1

# SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models

Abstract.
The rapid advancement of AI technologies, particularly Large Language Models (LLMs), is establishing a new paradigm for Business Intelligence (BI). Despite the emergence of pioneering work in enhancing BI systems with LLMs, we have identified the following three issues when deployed in real industrial scenarios: interaction limitations, performance bottlenecks, and functionality deficiencies.
In this paper, we present SiriusBI, an end-to-end business intelligence system that is designed to address the three issues simultaneously. First, we propose an intelligent and application-oriented module called multi-round dialogue with querying, which aims to overcome the prevalent interaction limitations in current BI solutions. Next, to mitigate the performance bottlenecks caused by scenario migration, we introduce two SQL generation methods that strike a balance between accuracy and deployment costs. Finally, to tackle the practical challenges posed by functionality deficiencies, we develop an end-to-end workflow that covers the entire BI process, ensuring that SiriusBI delivers a robust and complete set of functionalities.
As an independent cloud service in Tencent’s data platform, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients. Experiments on real-world datasets and practical applications in industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI. Remarkably, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
PVLDB Reference Format:
Jie Jiang,
Haining Xie,
Yu Shen,
Zihan Zhang,
Meng Lei,
Yifeng Zheng,
Yide Fang,
Chunyou Li,
Danqing Huang,
Wentao Zhang,
Yang Li,
Xiaofeng Yang,
Bin Cui,
Peng Chen
. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX
††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
Business Intelligence (BI) (Negash, 2004; Wieder and Ossimitz, 2015) is a crucial application scenario in the data field, comprising a comprehensive suite of methodologies, tools, and infrastructures designed to collect, integrate, analyze, and present an organization’s raw data to generate actionable insights for informed decision-making. The primary objective of BI is to generate actionable insights that facilitate informed decision-making. BI systems are extensively utilized across various sectors, including finance (Nuseir, 2021), environment (Hegger and Saraceni, 2024), and social media (Chen et al., 2024a; Rivera, 2024), which significantly enhances the decision-making process through the provision of real-time analytics and reporting capabilities.
A typical BI system comprises several key components: a data management system that stores, processes, and aggregates vast amounts of data; analytics algorithms that transform this data into actionable insights; and visualization tools that present the information in intuitive and user-friendly formats. These components collaboratively function to streamline business processes, enhance customer experiences, reduce operational costs, and mitigate risks. Recent advancements in LLMs have sparked significant interest in ChatBI (Lian et al., 2024) — an integrated BI system enhanced by conversational AI. Concurrently, the demand for a fully integrated and efficient ChatBI solution is surging, driven by the need of a more intuitive and accessible mode of data interaction. This evolution promises to transform how users engage with data, making insights more readily available and actionable.
To meet the growing demand for big data analysis and decision-making in BI, the data community has proposed numerous effective approaches. While NL2SQL is the core task in most BI systems, traditional methods often rely on the schema-based approaches (Hristidis and Papakonstantinou, 2002; Luo et al., 2007; Zeng et al., 2016; Hristidis et al., 2003) or parsing-based approaches (Iyer et al., 2017; Wang et al., 2017; Li and Jagadish, 2014; Popescu et al., 2004) to generate SQL queries from natural language. Recently, there has been a growing trend in implementing NL2SQL powered by LLMs, utilizing either prompt engineering (Pourreza and Rafiei, 2023; Wang et al., 2023) or model training techniques (Li et al., 2024; Pourreza and Rafiei, 2024). However, these methods typically focus on Single-Round Dialogue (SRD) scenarios, which limits their effectiveness in real-world applications where user interactions are typically more complex. To address the limitations of NL2SQL in BI contexts, the NL2BI (Lian et al., 2024) task has been introduced. This task aims to generate SQL queries based on Multi-Round Dialogues (MRD), allowing for a comprehensive understanding of user intentions. However, when implementing existing work in industrial BI systems and applying it in real-world scenarios, we identified the following three challenges:
C1: Interaction Limitations. ChatBI aims to leverage MRD to assist users in data analysis, thereby lowering the barrier to data usage and enhancing overall data utilization efficiency. However, existing public solutions still face challenges in supporting MRD due to real-world interaction modes. On one hand, much of the research (Wang et al., 2023; Talaei et al., 2024; Li et al., 2024; Pourreza and Rafiei, 2024, 2023) in the community concentrates on enhancing the performance of the NL2SQL task. While this has advanced the field to some extent, the interaction mode remains restricted to SRD, which imposes strict constraints on how users express their needs. This limitation ultimately undermines the primary objective of ChatBI systems: to facilitate universally accessible data analysis. On the other hand, leading technology teams in industry have introduced ChatBI (Lian et al., 2024) that support MRD and offered feasible methods for implementing multi-round interactions. Nevertheless, these solutions also impose limitations on user interaction modes in practical applications, catering only to fixed question-answer formats. Furthermore, they tend to be less effective in BI scenarios that require deep domain expertise.
C2: Performance Bottlenecks. The NL2SQL task is crucial for BI because it directly impacts the accuracy of data retrieval during analysis. In practical industrial scenarios, existing NL2SQL methods often exhibit suboptimal performance when applied across diverse contexts. For example, a LLM fine-tuned on data specifically annotated for the financial domain performs poorly when applied to the advertising domain. The process of preparing and standardizing data for each specialized field, followed by fine-tuning, incurs significant costs, which are often impractical in industrial applications.
C3: Functionality Deficiencies. Traditional BI systems are typically comprehensive solutions for data analysis, encompassing multiple interrelated modules such as data storage, SQL generation, and data insights. However, existing LLM-based BI systems fall short in providing essential functionalities, including efficient intent understanding, adaptive domain-specific knowledge, and intelligent data insights. For example, Mac-SQL (Wang et al., 2023) and CHESS (Talaei et al., 2024) primarily focus on NL2SQL tasks, while Lian et al. (2024) propose using the Apache SuperSet(Apa, [n.d.]) tool directly after NL2SQL, neglecting the components of domain knowledge management module and deep data insight. Unlike traditional BI systems, users expect LLMs to enhance the entire ChatBI workflow, which drives a technological paradigm shift that improves both application efficiency and user experience. A purely NL2SQL-based approach or scenario-constrained MRD functionality is inadequate to meet the real-world needs of BI applications. This gap presents an opportunity to develop a more comprehensive and intelligent ChatBI system.
To address the aforementioned challenges, we propose SiriusBI, which implements an end-to-end solution for BI scenarios. This system leverages the capabilities of LLMs to empower various submodules, thereby enhancing both efficiency and user experience in data analysis. Specifically, for the issue of interaction limitations (C1), we introduce the MRD-Q (Multi-Round Dialogue with Querying) module. Unlike the MRD functionality proposed by Lian et al. (2024), MRD-Q incorporates an intent clarification module that disambiguates user queries through follow-up questions. This approach enables the system to accurately identify the user’s true intent, even when the initial query is flawed or ambiguous, thereby facilitating precise responses, as shown in Figure 1. Furthermore, we employ a Retrieval-Augmented Generation (RAG) framework that leverages both business-domain knowledge and metadata, ensuring our system’s adaptability across various domains. Consequently, this reduces barriers for users and significantly enhances the overall user experience.
For the issue of performance bottlenecks (C2), we propose two implementation strategies in SQL generation. First, for domain-specific NL2SQL tasks, we introduce an efficient and automated process for constructing training data to improve the performance of LLMs. Second, for scenarios that require transferability, we present a two-step solution that integrates a Semantic Model with a SQL Model. This approach leverages domain-specific knowledge and metadata, ensuring high accuracy in SQL generation without the need of additional training. Furthermore, it reduces the costs associated with transferring between different scenarios and enhances the generalization of NL2SQL across various domains.
For the issue of functionality deficiencies (C3), SiriusBI provides a comprehensive end-to-end workflow. Unlike existing works that mainly focus on the NL2SQL task and the expansion of MRD modules, SiriusBI introduces two essential modules specifically designed for ChatBI : the Knowledge Management module and the Data Insights module. The knowledge management module is responsible for storing, processing, and utilizing domain-specific expertise and data to support the specialized knowledge requirements of various submodules. Meanwhile, the Data Insights module transcends traditional data visualization and report generation by providing critical key functionalities such as automatic data preparation, complex task planning, and the application of insight tools. These capabilities provide users with more comprehensive and in-depth support for their data analysis needs.
Contributions. The main contributions of this paper can be summarized as follows:
We propose SiriusBI, an end-to-end LLM-based ChatBI system designed to provide better data analysis services and lower the usage threshold for users. To the best of our knowledge, SiriusBI is the first comprehensive end-to-end ChatBI system that simultaneously addresses the aforementioned three challenges. So far, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients.
To extend the interaction modes of ChatBI systems, we introduce MRD-Q, a dialogue system designed to ask clarifying questions when user queries are unclear or ambiguous. This improves the accuracy of intent understanding, lowers the barrier to user entry, and ultimately enhances the overall user experience.
To enhance the transferability of ChatBI systems across different industrial scenarios, we propose two implementation approaches for SQL generation: single-step generation and two-step generation. The two-step generation method significantly improves the transferability of ChatBI and provides a new paradigm for SQL generation in the industry.
Extensive experiments on real-world datasets and industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI. Notably, when applied in industry, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
2. Preliminary
In this section, we introduce some key concepts including BI, ChatBI, NL2SQL, knowledge management and data insight.
2.1. BI and ChatBI
BI refers to the technologies, applications, and practices involved in the collection, integration, analysis, and presentation of business information. The primary goal of BI is to improve decision-making with business organizations. Traditional BI systems typically rely on data warehousing, data mining, and reporting tools, which enable organizations to transform raw data into meaningful insights.
ChatBI represent an evolution of traditional BI systems by integrating natural language processing (NLP) capabilities. This integration allows users to interact with BI tools using conversational interfaces. ChatBI systems aim to democratize data access by enabling non-technical users to query and analyze data through natural language queries. These systems leverage advanced NLP techniques to accurately understand user intent and generate relevant responses, thereby enhancing the overall user experience and allow organizations to make data-driven decisions more accessible and actionable.
2.2. NL2SQL Task
NL2SQL is a task in the field of natural language processing and database management that focuses on converting natural language queries into SQL queries.
In the specific task of SQL generation, we first receive the user’s query intent . Then, based on the relevant metrics and dimensions within the user’s intent, and in conjunction with the database schema information , and the SQL query result can be generated as,
where is the function that generates the SQL query result . The accuracy and efficiency of NL2SQL systems are critical for the effectiveness of ChatBI systems, as they directly impact the quality of the generated insights.
2.3. Knowledge Management
Knowledge Management (KM) involves systematic processes within an organization aimed at optimizing the use of knowledge. In most BI and ChatBI systems, KM is structured into three modules: knowledge processing, knowledge storage, and knowledge retrieval.
The critical formula for knowledge retrieval is:
where represents the structured knowledge from retrieval, represents the user’s query, denotes the stored knowledge, and is the function that facilitates the retrieval of knowledge from storage.
2.4. Data Insight
Data insights is a fundamental component of BI and ChatBI systems, designed to transform raw data into actionable insights through a series of structured processes: data aggregation, data analysis, insight generation, and visualization.
Data Aggregation: Combine data from multiple sources into a comprehensive dataset.
Data Analysis: Apply statistical and machine learning techniques to identify patterns and trends within the aggregated data.
Insight Generation: Translate the results of data analysis into meaningful insights that can inform business decisions.
Visualization: Present insights through intuitive visual formats to enhance understanding and facilitate decision-making.
3. System design
In this section, we introduce the overall architecture of SiriusBI. Compared with the current ChatBI systems (Lian et al., 2024), SiriusBI provides a more comprehensive workflow, especially in the knowledge management and data insight.
3.1. Overview
The overall framework of SiriusBI is illustrated in Figure 2 and consists of five key modules: (1) The Knowledge Management subsystem which stores, processes, and retrieves pertinent database knowledge, domain-specific expertise, and demonstrations (i.e., examples used for context learning in LLMs) that support the functionality of the other modules. (2) The Dialogue Analysis module, which accepts user queries and handles the interaction through multiple-round dialogues. (3) The Intention Querying module, which clarifies the user’s intent by generating follow-up questions based on the dialogue history. (4) The SQL Generation module, which converts the fully understood user intent into SQL queries. (5) The Data Insight module, which automatically prepares data, plans tasks, attributes issues, and predicts trends based on the data retrieved through NL2SQL queries.
3.2. Workflow
The workflow of SiriusBI is detailed as follows: When a user inputs a question in a BI scenario, for example, ”Company A’s year-to-date consumption compared year-over-year?” (), in the first stage, the Dialogue Analysis module evaluates whether the semantic content of the user’s input is complete, i.e., whether it includes both ”dimensions” and ”metrics”. If the input is determined to be semantically complete, it proceeds to the second stage, the Intention Querying module. However, if the user’s input is missing either dimensions or metrics, the module uses the conversation history to complete the semantics of the current question before proceeding to the next stage. For instance, if the user subsequently asks, ”How about Company B?” (), the Dialogue Analysis module will complete the question to ”Company B’s year-to-date expenditure compared year-over-year?” ().
In the second stage, the Intention Querying module recalls relevant information from the knowledge base based on the input question. Notably, to incorporate domain knowledge, improve system generalization, and enhance data insights, we have developed a specialized Knowledge Management module, which is elaborated in 4.1. After knowledge retrieval, we recall relevant fields from the database. If there is still ambiguity or confusion among the recalled fields, the Intention Querying module prompts the user for clarification to ensure accurate input for the SQL generation stage, thereby minimizing performance loss due to error propagation in the pipeline.
In the third stage, the SQL generation module receives the semantically completed user query from the previous stage and selects the appropriate tables for each query. Typically, this involves matching the top-k tables using database table information. To improve the adaptability of SiriusBI to various scenarios and requirements, we propose two optional generation methods: 1) An one-step generation method, where we construct specific datasets via an automated process for and perform Supervised Fine-Tuning (SFT) on tasks that require strong domain expertise but lower generalization; 2) A two-step SQL generation method, where we structure user queries via knowledge retrieval and then translate inputs into SQL using a fixed LLM. This method ensures fast deployment of SQL generation on unseen scenarios. By parallel use of these two generation methods, we maximize the accuracy of the generated SQL queries, ensuring reliable and relevant results for users.
In the fourth stage, the Data Insight module receives the SQL queries and their corresponding results from the third stage. Based on the user’s additional insight requirements, the module performs automatic data preparation, complex task planning, insight tool execution, and finally generates corresponding reports.
In the following section, we provide more in-depth exploration of the design and functionalities of each module.
4. System description
In this section, we explore various components of SiriusBI in detail. We begin with the Knowledge Management module in Section 4.1, which serves as a critical foundation that provides essential knowledge support for other parts of the workflow. Next, Section 4.2 focuses on Multi-round Dialogue with Querying, which consists of both the Dialogue Analysis and Intention querying module. In Section 4.3, we discuss SQL Generation, highlighting both the one-step and two-step SQL generation schemes. Finally, we introduce the Data Insight module in Section 4.4.
4.1. Knowledge Management
In industrial scenarios, the application of Business Intelligence (BI) often requires a substantial amount of domain-specific knowledge, including table metadata, business knowledge, and LLMs chat history, etc. To address the complex knowledge requirements, SiriusBI establishes the Knowledge Management module by constructing a multidimensional knowledge graph.
This knowledge graph comprises several key dimensions: 1) metadata information, which includes table schema, column values; 2) business knowledge, which includes domain-specific fields, term definitions, and relevant business context; 3) demonstration information utilized for prompting; 4) conversation data. Together, these dimensions establish a robust foundation for the knowledge management module.
The module also supports DSL (Domain-Specific Language) data generation, which includes functionalities such as query rewriting and DSL prompting. Furthermore, it facilitates API data retrieval and provides essential metadata and business knowledge, thereby supporting comprehensive data insights. This versatility makes the knowledge management module a vital component in various data-driven applications.
4.1.1. Knowledge Storage
For knowledge storage, we use EasyGraph (Gao et al., 2023) to construct a multidimensional knowledge graph. Figure 3 shows the graph stucture for storing knowledge related to data tables. Each node is a data structure that contains node types, descriptions, embeddings, and other information. The node types in this graph include databases, tables, columns, (column) values, UDFs (complex metric calculation functions such as month-over-month), terms (explanations of professional terms), and aliases (alternative names for other nodes). For columns, the node has specific attributes like whether it is a dimension or measure. The relationships between these nodes are represented as edges, forming a comprehensive knowledge storage structure used in SiriusBI. By defining graph structures and enhancing query efficiency through advanced graph algorithms, we not only organize knowledge effectively but also facilitate the rapid retrieval of relevant information during queries.
4.1.2. Knowledge Processing
Before storing the knowledge, it’s essential to process various types of data to ensure effective storage and management. We provide standard guidelines for the business teams to follow. To create a new node, domain experts first manually extract the node attributes based on specific business scenarios. Then, for knowledge with missing attributes, we complete it through metadata enhancement (Choudhury et al., 2023). This enhancement occurs on two levels: 1) automatic context completion using AI techniques, and 2) backfilling missing metadata through data engineering methods. Together, these processes ensure that the knowledge graph is ready for effective utilization.
4.1.3. Knowledge Retrieval
The knowledge retrieval process is divided into two main phases: coarse retrieval and fine retrieval. Figure 4 shows a brief overview of the retrieval process.
Coarse Retrieval. Initially, we use Elasticsearch (Elasticsearch, 2018) for coarse retrieval without segmenting the original query. Specifically, we deploy a hybrid search on both the lexical and semantic levels (Wang et al., 2024). To achieve a high recall rate, we allow for a larger recall set, temporarily prioritizing recall over precision. For each recalled node, if it is an alias or a value, we trace back to at least the column, term, or UDF level. For example, for SQL generation, this process yields all potentially relevant column names for fine retrieval.
Fine Retrieval. Based on the results from coarse retrieval, we perform a secondary search on the sub-graph obtained from the coarse retrieval by selecting the top-k nodes ranked by a reranking model (Chen et al., 2024b), which significantly reduces the number of candidate nodes. In addition to reranking, for SQL generation, we apply a LLM to filter out columns if it is not relevant to the user query.
4.2. Multi-Round Dialogue with Querying
Multi-Round Dialogue with Querying (MRD-Q) is a critical component in SiriusBI that receives and processes user input. It primarily supports multi-round dialogue (MRD), intent recognition, and clarification aimed at the user. This reduces the usage barrier, enhances the user experience, and minimizes the impact of errors caused by incorrect intent recognition. MRD-Q consists of two modules: dialogue analysis and intention querying. A detailed introduction is as follows:
4.2.1. Dialogue Analysis
In the industrial sector, BI has a wide range of applications aimed at enhancing user experience in data retrieval, data analysis, and data science. In this context, the multi-round dialogue feature in ChatBI plays a crucial role. However, most current NL2SQL research (Pourreza and Rafiei, 2023; Wang et al., 2023; Sun et al., 2023; Caferoğlu and Ulusoy, 2024; Li et al., 2024) is designed based on single-round dialogues, which creates a significant gap between this approach and real-world application scenarios. Firstly, single-round NL2SQL tasks require highly precise user input, meaning that even minor errors in the input may prevent the generation of correct SQL statements for subsequent data processing tasks. This high requirement not only severely diminishes user experience but also raises the barrier to entry for users who may not be familiar with SQL syntax. Secondly, data tasks, including data retrieval, are not isolated processes. Real-world data analysis or data science tasks often involve complex and detailed work, which inevitably requires extensive data retrieval efforts and even more sophisticated BI functionalities. The single-round dialogue approach for data retrieval is evidently inadequate to meet the demands of these real-world application scenarios.
The multi-round dialogue feature in SiriusBI is built upon the basic principles of dialogue analysis implementation in ChatBI(Lian et al., 2024). It determines the input for the SQL Generation module by assessing whether the user’s current input or historical conversation records contain indicators (columns) and dimensions. Specifically, the system first determines whether the user’s current input contains the two key pieces of information: indicators and dimensions. If both are found, the current input is considered semantically complete, and the workflow proceeds to the next module. If either is missing, the system supplements the current input’s semantics by selecting the most recent historical input that includes indicators and dimensions, and then moves to the next module.
Notably, unlike ChatBI (Lian et al., 2024), SiriusBI does not use traditional small NLP model fine-tuning to determine whether a statement contains indicators and dimensions or to select relevent historical conversations. Instead, we leverage the robust semantic understanding capabilities of LLMs along with the domain-specific knowledge in our knowledge base. This approach not only facilitates the system’s transition between different vertical domains but also enhances SiriusBI’s adaptability across various fields, which leads to a reduction on time and data construction costs associated with model training.
4.2.2. Intention Querying
To address the interaction limitations, we have added an intent clarification feature to the basic dialogue analysis module, resulting in a comprehensive MRD-Q component. The intention querying module is a critical supplement to dialogue analysis and includes the following three steps:
Intent Classification. When processing information from dialogue analysis, the first step is to reclassify the user’s true intentions. In multi-turn dialogues in real-world scenarios, certain metrics or dimensions may be missing, and relying solely on the dialogue history cannot complete the information. At this point, the results of intent classification are needed to decide whether to ask the user follow-up questions. We use the same LLM for semantic understanding as in Section 4.2.1 and categorize queries into three types. Denote the user input as , and the result of the intent classification as . We have,
When equals to 1, it indicates that, despite supplementing with historical inputs in Section 4.2.1, the user query still lacks critical information. Therefore, to ensure the successful execution of the task, we must either automatically retrieve essential information from the knowledge base or request additional details from the user.
Knowledge Retrieval. We utilize the knowledge retrieval function described in Section 4.1.3, using a hybrid approach of Embedding and ElasticSearch to retrieve demonstrations and domain knowledge. The demonstrations primarily serve as query clarification prompts. We have manually constructed a prompt library of common query clarification forms and stored it along with the necessary knowledge for SiriusBI in the knowledge management module.
Intent Clarification. Based on the results of intent recognition and knowledge retrieval, SiriusBI integrates and comprehensively analyzes these key pieces of information during the clarification phase. Specifically,
When is 0, the clarification phase will politely decline and guide the user to ask data analysis-related questions.
When is 1, the input is likely not reflected in the user’s historical information. In this case, the clarification phase will rely on the retrieved information from the knowledge base to enhance the indicators and dimensions. It will then ask the user if the supplemented parts meet their expectations. If not, the user is prompted to provide the relevant indicators and dimensions in the next round of dialogue. This iterative process helps clarify the user’s true intent through follow-up questions, i.e., the true dimensions and indicators of the query.
When is 2, no further clarification is needed, and the process directly moves to the next module in the SiriusBI pipeline.
By integrating the intent clarification feature into the dialogue analysis, we have developed the MRD-Q component. Through professionally designed clarification prompts, we can understand the user’s true intent more accurately. This approach lowers the user’s entry barrier. Even if the user provides arbitrary answers, SiriusBI can clarify the user’s correct intent through guided follow-up questions. Additionally, it helps to avoid confusion between similar dimensions and indicators, thereby preventing error propagation throughout the entire BI system.
4.3. SQL Generation
NL2SQL is the core task within the entire BI system, and the performance bottleneck continues to pose challenges in the industry. In this subsection, we will focus on the implementation approach of the SQL Generation module in SiriusBI, including table selection, as well as the SQL generation methods. To address the high cost of model migration across different professional domains, we have also proposed a novel two-step paradigm for SQL generation.
4.3.1. Table Selection
Before generating SQL queries, it is necessary to select the appropriate tables from the database based on the user’s input, especially retrieving the relevant table information mentioned in the user’s query. As shown in Figure 5, the implementation of table selection is divided into two parts: coarse ranking and reranking. In the coarse ranking stage, the number of candidate tables is reduced to a manageable number. In the reranking stage, these candidates are re-ranked to identify the most relevant ones for SQL generation.
The coarse ranking process consists of two parts: online and offline. In the offline part, each data table has a pre-prepared embedding. We use the StarRocks (SR) (team, [n.d.]) vector database to store the embeddings of tables to ensure data security. SR supports millisecond-level latency retrieval and allows for filtering by specified tags, making it a robust solution for efficient table selection. In the online part, we extract multiple keywords from the query through a LLM and then perform individual retrieval using each keyword. Compared to directly embedding the user’s input, this approach eliminates redundant information, reduces interference from irrelevant information, and increases the recall rate of target tables.
During reranking, we use different dimensional scoring methods for each table based on the hundreds of tables recalled in the coarse ranking stage. The overall calculation formula is as follows:
where represents the semantic recall score during the coarse ranking, which is the cosine similarity score of the vectorized results. represents the table’s popularity score, which is calculated by Tencent Data Warehouse based on query logs and execution task history. represents the similarity score between the user query and the table calculated from the field dimension, and , are hyperparameters.
Regarding the token-level scores, user queries frequently reference specific keywords, while tables may contain thousands of fields. Simple vectorizing the entire table’s information may obscure some information details, leading to suboptimal recall performance. To address this shortcoming, we design a token-level similarity calculation to achieve better granular information recall. Denote the i-th table as , and the j-th column in as . Assume that contains fields, and the input query contains keywords. Then, the corresponding similarity score of is the sum of similarity calculations between the table and keywords. The calculation formula is as follows,
where represents the similarity score between keyword and table , which is the maximum value of the text similarity with all columns:
We utilize TF-IDF (Joachims, 1997) as the vectorization method to compute similarity for several reasons: 1) The keywords are relatively short; 2) The calculation needs to be performed in real-time; 3) Gensim-based TF-IDF allows for rapid computation of similarity in batches; 4) Avoid the situation where the vectorized results fail to give proper attention to some keywords in the query when the input is excessively long.
4.3.2. One-Step SQL Generation
For scenarios where high accuracy in SQL generation is required and the usage domain is relatively stable, we apply a one-step approach that leverages supervised fine-tuning (SFT) of LLMs to allow direct generation of SQL statements from user inputs. Notably, we use Hunyuan-70B-SQL as the generation model, which is specifically designed for SQL generation by Tencent. While data quality is critical to the accuracy of the fine-tuned model, we develop an automated data pipeline which includes two critical steps: data preparation and data refinement.
Data Preparation. In the data preparation phase, we begin by selecting a batch of natural language requirements from business data, along with their ground-truth SQL generation results. Given that these requirements are highly standardized and contain complete contextual information, relying solely on this data poses challenges in ensuring the model’s performance in scenarios with ambiguous information.
To address this, we first employ a LLM to process each query, preserving the original intent while minimizing the information provided. This approach generates a set of semantically ambiguous inputs. In addition, we prepare relevant domain knowledge to serve as a reference for these inputs. We leverage knowledge retrieval techniques outlined in Section 4.1 to extract supplementary knowledge related to the original queries, while also enabling the LLM to supplement the rules it relies on during the information reduction process. An example of data preparation without ground-truth results is as below,
Finally, the dataset we use for fine-tuning comprises the reduced queries, the supplementary contextual knowledge, and the ground-truth SQL outputs. This comprehensive approach enhances the model’s ability to perform effectively in diverse and ambiguous scenarios.
Data Refinement. After fine-tuning on the prepared data, we further introduce an adaptive data refinement method to improve the accuracy of the generation model. We develop two strategies to improve the data quality based on error types and user feedbacks. The first strategy involves automatically assessing the model’s generated results and categorizing the failed examples into different types, such as time processing errors, unit conversion errors. Based on this information, we will prioritize queries that address these needs in the next data preparation phase. The second strategy involves conducting small-scale user testing, where users provide simple evaluations of the SQL queries generated by the model. Examples that receive low ratings will be annotated and added to the new training set. After evaluating the current model with these two strategies, we will iterate on the data preparation phase and fine-tune the model again. This process will continue until the model’s accuracy meets the requirements or until we reach the budget limit set by our clients.
4.3.3. Two-Step SQL Generation
To meet the demand for a method with strong generalization capabilities and low deployment costs across multiple SQL generation scenarios, we propose a novel two-step generation paradigm. This paradigm breaks down the SQL generation task into two stages: the Semantic Model stage and SQL Model stage.
In the Semantic Model stage, we first structure the user’s input. In practical NL2SQL tasks, user inputs are often difficult to control, and incomplete or ambiguous inputs can directly affect the SQL generation accuracy of ChatBI systems. This, in turn, reduces user experience and raises the user threshold. Therefore, we enhance the semantic information of the inputs before generating SQL from user inputs. For a user query , we first perform a retrieval from a knowledge base that includes explanations of domain-specific terms. Notably, our knowledge management module (Section 4.1) allows for plug-and-play management, which boosts efficiency of deployment in industrial applications. The retrieved domain knowledge, along with , is then processed by a LLM-based Semantic Model for semantic analysis and understanding, resulting in the generation of structured text. In our implementation, the structured text takes various forms, such as question rewriting, JSON, and API formats. Additionally, during the Semantic Model stage, we employ automated demonstration selection and Chain-of-Thought (CoT) prompting techniques. These strategies greatly enhance the accuracy and generalization capability of the Semantic Model for different types of questions.
In the SQL Model stage, we use another LLM as the SQL generation model, specifically Hunyuan-70B-SQL. The enhanced structured information from the Semantic Model serves as the input for the final SQL generation. As the structured input provides ample context for SQL generation, the SQL Model can effortlessly generate SQL dialects tailored to various database systems from the same input. By structuring and supplementing user questions before generating SQL, the two-step generation paradigm ensures the accuracy of SQL generation, and meanwhile facilitates easy migration across different industrial scenarios. This approach significantly reduces the training costs and improves practical deployment efficiency.
| Category | Example | Description |
| Descriptive Requests | Which day had the highest revenue? | This category is a general term for basic analysis problems, including correlation, extreme value, difference analysis, etc. |
| Predictive Requests | Forecast failure rate for the next week. | This category include future trend analysis or label prediction of unseen data. |
| Diagnostic Requests | Identify anomalies in the sales revenue. | This category includes anomaly detection and metric attribution for numerical data, as well as drill-down for categorical data. |
| Prescriptive Requests | How to deal with abnormal values of production volume? | This category helps interpret data insights and proposes actionable solutions based on the analysis. |
| Complex Requests | Forecast the net income for next week and calculate the expected year-over-year change. | This category involves executing multiple sub-tasks and there are dependencies between the execution of these sub-tasks. |
4.4. Data Insight
Data insights are essential to ChatBI systems, as they enable users to extract valuable information from raw data. This functionality allows users not only to understand the current state of affairs but also forecast future trends. Data insights go beyond traditional descriptive analytics; they reveal hidden patterns in the data and provide deep understanding, including root cause analysis of issues, predictive modeling for the future, and actionable recommendations based on the analysis. Such in-depth analysis is crucial for businesses because it helps decision-makers make informed choices, thus driving business growth. SiriusBI leverages the capabilities of LLMs to refine the functionalities of data insights, which greatly improves the user experience. The data insight module of SiriusBI is designed to address various problems as detailed in Table 1.
4.4.1. Main Workflow
The primary workflow for data insights in SiriusBI includes the following steps: automatic data preparation, complex task planning, invocation of insight tools, and generation of insight reports. A more detailed process flow will be introduced as follows.
Automatic Data Preparation. The insight tool executes three sequential steps during automatic data preparation: intent rewriting, Text2DSL, and metric-dimension recommendation. Initially, intent rewriting categorizes and refines the given insight question into a more specific description. Subsequently, Text2DSL converts the question into a Domain-Specific Language (DSL) corresponding to SQL queries. By inputting the schema information of the data table along with the insight question, this part outputs structured DSL language that includes essential columns for data insight. Finally, the schema of the data table, the DSL output, and the insight question are input into the metric-dimension recommendation part, which suggests additional relevant data columns that could fulfill the insight requirements based on the type of insight task.
Complex Task Planning. The Data Insight module is capable of planning complex tasks, in which a question may contains multiple interdependent requirements. It employs an internal planner that decompose these tasks into manageable sub-tasks and iteratively refines the plan based on the execution results of the previous sub-task. For instance, consider a scenario where the module is tasked with calculating the month-over-month growth of net new holdings for July 8, 2024, using a business dataset. Since the data for that specific date is not available in the data table, the execution process must first include a predictive analysis to estimate the value for July 8 before proceeding with the growth calculation.
Insight Tool Execution. The capabilities of the data insight module are supported by specialized insight tools. Depending on the specific tasks, the insight tools in SiriusBI primarily include:
Code Generation. For straightforward queries, such as retrieving a specific piece of data, SiriusBI utilizes a LLM to generate the necessary code, which is then executed in a secure sandbox environment.
Data Modeling. For specialized tasks such as prediction and anomaly detection, SiriusBI develops specialized machine learning models and utilizes automated tools for hyperparameter optimization, providing users with accurate output support.
Visualization. For visualization needs, SiriusBI uses a LLM to construct configuration files of visualization settings, which are then used by underlying visualization tools to automatically generate charts.
Knowledge Retrieval. For questions that go beyond the information in the data tables, SiriusBI employs a knowledge retrieval function to extract relevant information (e.g., actions taken in similar scenarios) from the knowledge management module introduced in Section 4.1, thereby assisting in addressing the current task.
Insight Summary. Insight summary is a comprehensive tool to meet various insight requirements. It automatically generates data overviews, trend analyses, anomaly diagnostics, and summary recommendations based on the characteristics of the given data table as well as the outputs from previous insight tools.
Compared with traditional BI systems and the most advanced ChatBI systems, the data insight module proposed in SiriusBI offers users enhanced data analysis capabilities. This module supports a wide range of practical application scenarios, enabling users to achieve more usable and reliable data analysis results. In addition, it addresses complex and challenging data analysis tasks, meeting the growing demand for dependable analytical solutions.
5. Experiments and Results
To evaluate our framework, in this section, we conduct an in-depth investigation from two insights: 1) On the general NL2SQL task, SiriusBI outperforms the state-of-the-art baselines on both SRD and MRD datasets; 2) The end-to-end design of the SiriusBI system is rational, where the individual performance of each module demonstrates significant superiority.
5.1. Experiment Setups
Datasets. To demonstrate the overall performance of SiriusBI and the effectiveness of its various modules, we conduct comprehensive experiments using datasets from multiple tasks. Detailed descriptions of the datasets used for each task are provided below:
NL2SQL. Although the Spider (Yu et al., 2018) and Bird (Li et al., 2023) datasets are widely used for NL2SQL tasks, their default SRD settings and the gap with industrial scenarios make it difficult to accurately evaluate NL2SQL tasks in actual BI scenarios. Therefore, we follow the previous approach (Lian et al., 2024) and select high-quality and reliable data from real application scenarios to construct a muilt-round dialogue (MRD) dataset. The dataset contains 55 sets of queries, where each set consists of 2 to 5 rounds. In total, the dataset contains 160 queries spanning 35 columns, accompanied by SQL written by human experts for reference. Among these queries, 10 sets do not fully include all metrics and dimensions, while another 10 sets exhibit confusion in metrics or dimensions. While most baselines are proposed for single-round dialogue (SRD) scenario, to ensure fair comparison, we also supplement the missing information in the MRD dataset and add 22 more single-round queries, thereby creating a SRD dataset with totally 182 queries for evaluation purposes.
Knowledge Management. To evaluate the knowledge management module, we focus on two main aspects: 1) Assessing its impact on NL2SQL tasks; 2) Evaluating the performance of knowledge retrieval. In the latter evaluation, we utilize the knowledge base corresponding to the MRD dataset, which contains knowledge across six dimensions: table, column, value, term, udf, and alias, totaling 7704 pieces of knowledge.
Data Insight. For evaluation of data insights, we select 35 descriptive requests and 45 complex requests involving descriptive analysis used in industrial scenarios. The requests involve data retrieval, extreme value calculation, correlation analysis, etc. Note that, other types of requests are not included in the dataset, as it’s difficult to evaluate the correctness of answers to open-ended questions.
Evaluation Metrics. In NL2SQL tasks, we follow previous work (Lian et al., 2024) and applies the Useful Execution Accuracy (UEX) metric. While other metrics like Execution Accuracy and Exact Match require the generated query matches the ground-truth query precisely, UEX recognizes a query as correct if the execution results of the generated query align with the actual results.
For tasks beyond NL2SQL within SiriusBI, such as knowledge retrieval in knowledge management, comprehensive analysis in data insight, we employ the retrieval recall rate and the execution accuracy rate, respectively.
Baselines. We compare the NL2SQL capability of SiriusBI with three state-of-the-art baselines as below:
DIN-SQL (Pourreza and Rafiei, 2023) breaks down the NL2SQL task into discrete sub-tasks, providing specific prompts for each sub-task to GPT-4o through the completion of each component.
MAC-SQL (Wang et al., 2023) employs a multi-agent framework to handle NL2SQL tasks, involving three main agents: selector, decomposer, and refiner. These agents work together to enhance the accuracy and efficiency of SQL generation.
MRD-SQL is a reproduction of previous work (Lian et al., 2024). We reproduce the basic MRD system integrated with an SQL generation model. To ensure fair comparison, we use the same base LLM as in SiriusBI.
SiriusBI employs all modules introduced in this paper. For evaluation of NL2SQL, the data insight module is excluded. SiriusBI processes user inputs through dialogue analysis and intention querying before generating SQL queries. We use “one-step” and “two-step” to distinguish between the different SQL generation methods in SiriusBI.
LLM Settings. In SiriusBI and MRD-SQL, we utilize the Hunyuan-70B-SQL, specifically designed for SQL scenarios, developed by Tencent’s Hunyuan LLM group. For other baselines, we employ the most advanced OpenAI API (Ope, [n.d.]) as the foundational LLM, specifically the ”gpt-4o-2024-08-06” version, to ensure optimal performance in all compared baselines.
Training Setting. For LLM fine-tuning in one-step SQL generation method, we use four nodes equipped with a total of 32 NVIDIA-A100-80GB GPUs, setting the training to 3 epochs, a learning rate of 3e-5, a batch size of 2, and a sequence length of 8192 tokens. The fine-tuning is conducted on a private dataset consisting of over 9,000 entries collected from industrial scenarios.
5.2. Experiment with NL2SQL
In this experiment, we primarily evaluate the performance of various baselines on the SRD and MRD datasets. Note that, as an analysis module following NL2SQL, the data insights module do not participate in this part.
5.2.1. Results on SRD Dataset
We first conduct experiments on SRD dataset and present the results in Table 2. Our findings indicate that SiriusBI with both two SQL generation methods show a clear improvement over other baselines on the dataset. Specifically, when compared to competitive baselines such as GPT-4o and its derivatives (DIN-SQL and MAC-SQL), SiriusBI achieves an improvement of at least 40% in UEX. In comparison to the SOTA ChatBI implementation, MRD-SQL, the two-step and one-step variants of SiriusBI achieve improvements of 17.34% and 21.39%, respectively, which highlights the effectiveness of SiriusBI NL2SQL pipeline.
Interestingly, there is a relatively slight performance difference between two variants of SiriusBI. Concretely, the one-step method, which has been carefully fine-tuned on domain-specific data, gets 4.04% higher in UEX than the SFT-free two-step method. However, it is worth noting that the data for SFT took two weeks of annotation by four experts, and SFT itself also took 12 hours. Therefore, we recommend that users seeking higher accuracy use the one-step approach, while those prioritizing faster deployment efficiency should opt for the two-step approach.
| Baseline | UEX |
| DIN-SQL | 30.06% |
| MAC-SQL | 36.42% |
| MRD-SQL | 61.85% |
| SiriusBI two-step | 79.19% |
| SiriusBI one-step | 83.24% |
5.2.2. Results on MRD Dataset
Next, we evaluate SiriusBI on the MRD dataset and present the specific experimental results in Table 4. Our findings reveal that single-round methods perform poorly on the MRD dataset, where the UEX falls below 16%. In contrast, the two-step and one-step variants of SiriusBI achieve UEX scores of 55.00% and 62.50%, respectively, demonstrating the superiority of SiriusBI in real-world NL2SQL tasks when compared to the state-of-the-art multi-round design, MRD-SQL.
Remarkably, we also investigate the impact of SFT on the MRD dataset. By comparing the performance of SiriusBI one-step and MRD-SQL with and without fine-tuning, we observe that SFT yields significant performance gains at the cost of higher resource investment. Furthermore, we find that though SiriusBI two-step does not incorporate SFT, it still improves SiriusBI one-step without SFT by 5.62% through effective retrieval of domain knowledge. This leads to a similar conclusion that users can choose the one-step method for higher accuracy and the two-step method for faster deployment.
5.2.3. In-production Results.
Besides experiments on datasets collected from industrial scenarios, we also evaluate SiriusBI in production. So far, SiriusBI has been deployed in Tencent’s finance, advertising and cloud sectors. For Tencent Finance and Advertisement, we apply the one-step method by fine-tuning LLMs on elaborately collected domain-specific datasets, achieving in-production accuracies are 97% and 89%, respectively. In the case of Tencent Cloud where SiriusBI is available to business clients, we apply the two-step method to reduce deployment costs for users. The average accuracy across 14 clients are 91%. Interestingly, we find that the in-production results are higher than those reported on the SRD and MRD datasets. We attribute this to the differences in problem distribution. In industrial scenarios, users’ questions may be simpler, while the datasets include high-quality and complex questions, which are more suitable for assessing the system’s specialized capabilities.
| Baseline | UEX |
| DIN-SQL | 11.25% |
| MAC-SQL | 15.63% |
| MRD-SQL w/o SFT | 22.50% |
| MRD-SQL | 28.13% |
| SiriusBI two-step | 55.00% |
| SiriusBI one-step w/o SFT | 49.38% |
| SiriusBI one-step | 62.50% |
5.3. Systematic Deep Analysis Experiments
In addition to NL2SQL, we also conduct a deeper analysis of the effectiveness of various components in SiriusBI. In this section, we primarily focus on analyzing modules related with MRD-Q, knowledge management and data insights.
5.3.1. Evaluation of MRD-Q
In this part, we present the ablation study on dialogue analysis and intention querying using the MRD dataset/ The experimental results are displayed in Table 4.
Based on the experimental results, we can observe that, both the SiriusBI one-step and SiriusBI two-step heavily rely on the dialogue analysis and the intention querying modules. Specifically, when both modules are removed (without MRD-Q), the performance of SiriusBI declines significantly, with drops of 34.37% and 35.62% for one-step and two-step designs, respectively. Similarly, the absence of the intention querying module results in a performance degradation of 11.25% and 21.25% for SiriusBI. This indicates that the dialogue analysis and intention querying modules are critical to the overall effectiveness of SiriusBI.
In addition, the ablation study reveals that SiriusBI one-step fine-tuned on a private dataset exhibits greater robustness compared to SiriusBI two-step. Concretely, the impact of removing intention querying on SiriusBI one-step is less pronounced, resulting in a performance drop of only 11.25%, compared to a more significant decline of 21.25% for SiriusBI two-step. The difference can be attributed to the fact that SiriusBI one-step acquires domain-specific knowledge from the training data during the learning process, which helps to mitigate the negative effects of lacking conversational understanding and intention querying.
| Baseline | UEX | |
| SiriusBI two-step w/o MRD-Q | 20.62% | -34.38% |
| SiriusBI two-step w/o Q | 33.75% | -21.25% |
| SiriusBI two-step | 55.00% | - |
| SiriusBI one-step w/o MRD-Q | 46.88% | -15.62% |
| SiriusBI one-step w/o Q | 56.25% | -6.25% |
| SiriusBI one-step | 62.50% | - |
5.3.2. Evaluation of Knowledge Management
Then, we examine the impact of knowledge management (KM) on the NL2SQL pipeline. The results of disabling this module on the MRD dataset are presented in Table 5. The findings reveal a significant decline in performance for SiriusBI when KM is turned off. Concretely, UEX decreases by 23.12% for SiriusBI one-step and 26.87% for SiriusBI two-step. These results highlight the critical role of the knowledge management module in enhancing the overall performance of NL2SQL, as both intention querying and SQL generation heavily depend on the availability of relevant knowledge.
| Baseline | UEX | |
| SiriusBI two-step w/o KM | 28.13% | -26.87% |
| SiriusBI two-step | 55.00% | - |
| SiriusBI one-step w/o KM | 39.38% | -23.12% |
| SiriusBI one-step | 62.50% | - |
Next, we evaluate the module using the dataset for knowledge retrieval with 7704 samples. In this evaluation, we compare SiriusBI against widely used baseline models BM25 (Robertson and Sparck Jones, 1976) and Vector (Karpukhin et al., 2020), both of which are prominent in the field of text retrieval. We assess the performance of these methods using the recall rates of knowledge retrieval for all keywords (KR) and schema linking (SL) for table columns only. Our results indicate that SiriusBI outperforms the baselines across both metrics, thereby ensuring that other modules leveraging knowledge management within SiriusBI can access accurate and relevant additional information.
| Baseline | KR Recall | SL Recall |
| BM25 | 90.28% | 89.58% |
| Vector | 93.06% | 93.06% |
| SiriusBI | 95.14% | 93.75% |
5.3.3. Data Insight Evaluations
Finally, we conduct experiments on the Data Insight module. Remind that the data insight dataset contains quantifiable questions involving comprehensive analysis and complex task planning.
| Baseline | Accuracy |
| Command-R+ | 75% |
| Qwen2.5 | 80% |
| GPT-4o | 83% |
| SiriusBI | 95% |
For Comprehensive Analysis tasks, we use 100 real-world queries from the financial domain and conducts comparative experiments with three powerful LLM baselines: GPT-4o (Ope, [n.d.]), Qwen2.5 (Team, 2024) and Commnd R+ (cohere.ai Team, [n.d.]). The accuracy of responses to these queries serves as the evaluation criterion, and to ensure the reliability of the experimental results, each query is tested for three times. The specific experimental results are shown in Table 7. The findings indicate that the Data Insight module of SiriusBI maintains a notable performance advantage even against the currently most effective LLMs. The function call architecture (Kim et al., 2024) in SiriusBI enables it to outperform GPT-4o by 12% in Comprehensive Analysis tasks.
For Complex Task Planning, we compare the Data Insight module with a method that only performs tool invocation on 50 high-quality queries. Compared to the simple tool invocation method, the multi-agent architecture adopted by SiriusBI shows a significant advantage, by improving the accuracy from 70% to 100% on this dataset. To facilitate understanding of the execution process of Complex Task Planning, we provide an example as follows:
6. Related Work
The integration of Artificial Intelligence (AI) and Machine Learning (ML) has fundamentally transformed data-driven decision-making and strategic planning within the realm of Business Intelligence (BI) (Rane et al., 2024). Machine Learning and Predictive Analytics are essential for enhancing BI, providing advanced tools for forecasting and identifying patterns in extensive datasets (Lee et al., 2022; Tamang et al., 2021). Additionally, Natural Language Processing (NLP) and Text Analytics enable the extraction of valuable insights from unstructured data, empowering businesses to understand customer sentiment, automate customer service, and enhance decision-making processes (Kaur et al., 2019; Ittoo et al., 2016). By leveraging the Internet of Things (IoT) and Big Data, businesses can increase their resilience and adaptability in an ever-evolving market (Alghamdi and Al-Baity, 2022; Klisarova-Belcheva et al., 2017). Advanced AI models, such as ChatGPT (Ope, [n.d.]), play a significantly augment BI by providing intelligent data interpretations and improving human-machine interactions, making data more accessible and comprehensible for non-technical users (Linkon et al., 2024).
A BI system typically stores vast amounts of information, such as customer and product sales data. In traditional BI systems, users need to use SQL to aggregate the required data and perform downstream analysis. Therefore, to lower the barrier to using BI capabilities and to create an intelligent BI system, NL2SQL has become a core task. Recently, Large Language Models (LLMs), including ChatGPT (Ope, [n.d.]), Codex (cod, [n.d.]) and Claude 3 (cla, [n.d.]), have demonstrated remarkable potential in addressing NL2SQL tasks. These advancements benefit from the introduction of prompt engineering (Giray, 2023), in-context learning (Dong et al., 2022), and the chain-of-thought approach (Wei et al., 2022). Research into prompt-based NL2SQL methods has not only provided new technical paradigms for addressing the challenges of this task but has also revealed opportunities for reducing technical costs in industrial applications. For instance, methods such as DIN-SQL (Pourreza and Rafiei, 2023), C3 (Dong et al., 2023), MAC-SQL (Wang et al., 2023), and SQL-PaLM (Sun et al., 2023), designed for models like ChatGPT and PaLM (Narang and Chowdhery, 2022), have demonstrated significant improvements in the performance on NL2SQL tasks.
Prior to the emergence of LLMs, solutions to the NL2SQL problem typically involved training an ”encoder-decoder” model in a sequence-to-sequence format (Cai et al., 2021; Cao et al., 2021; Choi et al., 2021; Gan et al., 2021; Wang et al., 2020), which heavily relied on a substantial amount of database and SQL-related data. The rise of large models has introduced new technical paradigms for traditional train-based methods. For example, Codes (Li et al., 2024) propose a training method tailored for the NL2SQL task, which enhances the performance of LLMs in this area. However, to achieve better performance in the NL2SQL task, some studies (Maamari et al., 2024; Pourreza et al., 2024; Caferoğlu and Ulusoy, 2024) opt to train LLMs such as ChatGPT (Ope, [n.d.]) and Gemini (gem, [n.d.]), which often comes with a high cost in practical applications.
In summary, although there have been pioneering efforts in academia to integrate LLMs into BI systems, these works have not fully addressed practical needs. From a practical perspective, we must also consider the flexible user interaction requirements, efficient NL2SQL capabilities, and a complete BI workflow. Therefore, how to incorporate LLMs into BI and develop a new generation of intelligent BI tools remains an open question.
7. Conclusion
In this paper, we propose SiriusBI, an deployed end-to-end LLM-based Business Intelligence system specifically designed to address three key challenges encountered in industrial scenarios. Our approach includes: 1) implementing multi-round dialogue with querying capabilities to overcome interaction limitations; 2) employing two SQL generation strategies to alleviate performance bottlenecks; and 3) establishing a comprehensive end-to-end workflow of the entire BI process to address functionality deficiencies. Currently, SiriusBI is deployed across Tencent’s finance, advertising and cloud sectors, serving dozens of enterprise clients. Extensive experiments on industrial datasets demonstrate that SiriusBI clearly outperforms competitive LLM-based baselines in both SQL generation and data insights generation.
Acknowledgements.
We thank the researchers and engineers from the Data Platform department, Technology and Engineering Group (TEG) of Tencent for their technical and engineering support.References
- (1)
- Apa ([n.d.]) [n.d.]. Apache Superset. https://superset.apache.org/
- cla ([n.d.]) [n.d.]. Claude3. https://anthropic.com/claude
- gem ([n.d.]) [n.d.]. Gemini. https://gemini.google.com/
- Ope ([n.d.]) [n.d.]. Introduction of OpenAI Text Generation APIs. https://platform.openai.com/docs/guides/text-generation
- cod ([n.d.]) [n.d.]. OpenAI CodeX. https://openai.com/index/openai-codex
- Alghamdi and Al-Baity (2022) Noorah A. Alghamdi and Heyam H. Al-Baity. 2022. Augmented Analytics Driven by AI: A Digital Transformation beyond Business Intelligence. Sensors 22, 20 (2022), 8071.
- Caferoğlu and Ulusoy (2024) Hasan Alp Caferoğlu and Özgür Ulusoy. 2024. E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL. arXiv preprint arXiv:2409.16751 (2024).
- Cai et al. (2021) Ruichu Cai, Jinjie Yuan, Boyan Xu, and Zhifeng Hao. 2021. Sadga: Structure-aware dual graph aggregation network for text-to-sql. Advances in Neural Information Processing Systems 34 (2021), 7664–7676.
- Cao et al. (2021) Ruisheng Cao, Lu Chen, Zhi Chen, Yanbin Zhao, Su Zhu, and Kai Yu. 2021. LGESQL: Line Graph Enhanced Text-to-SQL Model with Mixed Local and Non-Local Relations. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing. 2541–2555.
- Chen et al. (2024b) Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian, and Zheng Liu. 2024b. Bge m3-embedding: Multi-lingual, multi-functionality, multi-granularity text embeddings through self-knowledge distillation. arXiv preprint arXiv:2402.03216 (2024).
- Chen et al. (2024a) Xiaoxin Chen, Meng Wu, and Mangning Wang. 2024a. Application of business intelligence under deep neural network in credit scoring of bank users. Journal of Computational Methods in Sciences and Engineering 24, 3 (2024), 1585–1604.
- Choi et al. (2021) DongHyun Choi, Myeong Cheol Shin, EungGyun Kim, and Dong Ryeol Shin. 2021. Ryansql: Recursively applying sketch-based slot fillings for complex text-to-sql in cross-domain databases. Computational Linguistics 47, 2 (2021), 309–332.
- Choudhury et al. (2023) Muntabir Hasan Choudhury, Lamia Salsabil, Himarsha R Jayanetti, Jian Wu, William A Ingram, and Edward A Fox. 2023. MetaEnhance: Metadata Quality Improvement for Electronic Theses and Dissertations of University Libraries. In 2023 ACM/IEEE Joint Conference on Digital Libraries. 61–65.
- cohere.ai Team ([n.d.]) cohere.ai Team. [n.d.]. The Cohere Platform. https://docs.cohere.com
- Dong et al. (2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Jingyuan Ma, Rui Li, Heming Xia, Jingjing Xu, Zhiyong Wu, Tianyu Liu, et al. 2022. A survey on in-context learning. arXiv preprint arXiv:2301.00234 (2022).
- Dong et al. (2023) Xuemei Dong, Chao Zhang, Yuhang Ge, Yuren Mao, Yunjun Gao, Jinshu Lin, Dongfang Lou, et al. 2023. C3: Zero-shot text-to-sql with chatgpt. arXiv preprint arXiv:2307.07306 (2023).
- Elasticsearch (2018) BV Elasticsearch. 2018. Elasticsearch. software], version 6, 1 (2018).
- Gan et al. (2021) Yujian Gan, Xinyun Chen, Jinxia Xie, Matthew Purver, John R Woodward, John Drake, and Qiaofu Zhang. 2021. Natural SQL: Making SQL Easier to Infer from Natural Language Specifications. arXiv preprint arXiv:2109.05153 (2021).
- Gao et al. (2023) Min Gao, Zheng Li, Ruichen Li, Chenhao Cui, Xinyuan Chen, Bodian Ye, Yupeng Li, Weiwei Gu, Qingyuan Gong, Xin Wang, et al. 2023. EasyGraph: A multifunctional, cross-platform, and effective library for interdisciplinary network analysis. Patterns 4, 10 (2023), 100839.
- Giray (2023) Louie Giray. 2023. Prompt engineering with ChatGPT: a guide for academic writers. Annals of biomedical engineering 51, 12 (2023), 2629–2633.
- Hegger and Saraceni (2024) Marco Hegger and Adriana Saraceni. 2024. Smart Port Sustainability: A Business Intelligence Framework for CO2 Reduction in Cargo Truck Operations. In Advances in Production Management Systems. Production Management Systems for Volatile, Uncertain, Complex, and Ambiguous Environments - 43rd IFIP WG 5.7 International Conference, Vol. 728. 309–323.
- Hristidis et al. (2003) Vagelis Hristidis, Luis Gravano, and Yannis Papakonstantinou. 2003. Efficient IR-Style Keyword Search over Relational Databases. In Proceedings of 29th International Conference on Very Large Data Bases. 850–861.
- Hristidis and Papakonstantinou (2002) Vagelis Hristidis and Yannis Papakonstantinou. 2002. DISCOVER: Keyword Search in Relational Databases. In Proceedings of 28th International Conference on Very Large Data Bases. 670–681.
- Ittoo et al. (2016) Ashwin Ittoo, Antal van den Bosch, et al. 2016. Text analytics in industry: Challenges, desiderata and trends. Computers in Industry 78 (2016), 96–107.
- Iyer et al. (2017) Srinivasan Iyer, Ioannis Konstas, Alvin Cheung, Jayant Krishnamurthy, and Luke Zettlemoyer. 2017. Learning a Neural Semantic Parser from User Feedback. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics. 963–973.
- Joachims (1997) Thorsten Joachims. 1997. A Probabilistic Analysis of the Rocchio Algorithm with TFIDF for Text Categorization. In Proceedings of the Fourteenth International Conference on Machine Learning. 143–151.
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick S. H. Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing. 6769–6781.
- Kaur et al. (2019) Jagbir Kaur, Ashok Choppadandi, Pradeep Kumar Chenchala, Varun Nakra, and Pandi Kirupa Gopalakrishna Pandian. 2019. Case Studies on Improving User Interaction and Satisfaction using AI-Enabled Chatbots for Customer Service. International Journal of Transcontinental Discoveries 6, 1 (2019), 29–34.
- Kim et al. (2024) Sehoon Kim, Suhong Moon, Ryan Tabrizi, Nicholas Lee, Michael W Mahoney, Kurt Keutzer, and Amir Gholami. 2024. An LLM Compiler for Parallel Function Calling. In Forty-first International Conference on Machine Learning.
- Klisarova-Belcheva et al. (2017) S Klisarova-Belcheva, G Ilieva, and T Yankova. 2017. BUSINESS INTELLIGENCE AND ANALYTICS–CONTEMPORARY SYSTEM MODEL. Trakia Journal of Sciences 15, 1 (2017), 298–304.
- Lee et al. (2022) Chee Sun Lee, Peck Yeng Sharon Cheang, and Massoud Moslehpour. 2022. Predictive analytics in business analytics: decision tree. Advances in Decision Sciences 26, 1 (2022), 1–29.
- Li and Jagadish (2014) Fei Li and H. V. Jagadish. 2014. Constructing an Interactive Natural Language Interface for Relational Databases. Proceedings of the Very Large Data Bases Endowment 8, 1 (2014), 73–84.
- Li et al. (2024) Haoyang Li, Jing Zhang, Hanbing Liu, Ju Fan, Xiaokang Zhang, Jun Zhu, Renjie Wei, Hongyan Pan, Cuiping Li, and Hong Chen. 2024. CodeS: Towards Building Open-source Language Models for Text-to-SQL. Proceedings of the ACM on Management of Data 2, 3 (2024), 1–28.
- Li et al. (2023) Jinyang Li, Binyuan Hui, Ge Qu, Jiaxi Yang, Binhua Li, Bowen Li, Bailin Wang, Bowen Qin, Ruiying Geng, Nan Huo, Xuanhe Zhou, Chenhao Ma, Guoliang Li, Kevin Chen-Chuan Chang, Fei Huang, Reynold Cheng, and Yongbin Li. 2023. Can LLM Already Serve as A Database Interface? A BIg Bench for Large-Scale Database Grounded Text-to-SQLs. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023.
- Lian et al. (2024) Jinqing Lian, Xinyi Liu, Yingxia Shao, Yang Dong, Ming Wang, Zhang Wei, Tianqi Wan, Ming Dong, and Hailin Yan. 2024. ChatBI: Towards Natural Language to Complex Business Intelligence SQL. arXiv preprint arXiv:2405.00527 (2024).
- Linkon et al. (2024) Ahmed Ali Linkon, Mujiba Shaima, Md Shohail Uddin Sarker, Norun Nabi, Md Nasir Uddin Rana, Sandip Kumar Ghosh, Mohammad Anisur Rahman, Hammed Esa, Faiaz Rahat Chowdhury, et al. 2024. Advancements and applications of generative artificial intelligence and large language models on business management: A comprehensive review. Journal of Computer Science and Technology Studies 6, 1 (2024), 225–232.
- Luo et al. (2007) Yi Luo, Xuemin Lin, Wei Wang, and Xiaofang Zhou. 2007. Spark: top-k keyword query in relational databases. In Proceedings of the International Conference on Management of Data. ACM, 115–126.
- Maamari et al. (2024) Karime Maamari, Fadhil Abubaker, Daniel Jaroslawicz, and Amine Mhedhbi. 2024. The Death of Schema Linking? Text-to-SQL in the Age of Well-Reasoned Language Models. In NeurIPS 2024 Third Table Representation Learning Workshop.
- Narang and Chowdhery (2022) Sharan Narang and Aakanksha Chowdhery. 2022. Pathways language model (palm): Scaling to 540 billion parameters for breakthrough performance. Google AI Blog (2022).
- Negash (2004) Solomon Negash. 2004. Business Intelligence. Communications of the Association for Information Systems 177, 195 (2004), 177.
- Nuseir (2021) Mohammed T Nuseir. 2021. Designing business intelligence (BI) for production, distribution and customer services: a case study of a UAE-based organization. Business Process Management Journal 27, 4 (2021), 1275–1295.
- Popescu et al. (2004) Ana-Maria Popescu, Alex Armanasu, Oren Etzioni, David Ko, and Alexander Yates. 2004. Modern Natural Language Interfaces to Databases: Composing Statistical Parsing with Semantic Tractability. In Proceedings of the 20th international conference on Computational Linguistics.
- Pourreza et al. (2024) Mohammadreza Pourreza, Hailong Li, Ruoxi Sun, Yeounoh Chung, Shayan Talaei, Gaurav Tarlok Kakkar, Yu Gan, Amin Saberi, Fatma Ozcan, and Sercan O Arik. 2024. CHASE-SQL: Multi-Path Reasoning and Preference Optimized Candidate Selection in Text-to-SQL. arXiv preprint arXiv:2410.01943 (2024).
- Pourreza and Rafiei (2023) Mohammadreza Pourreza and Davood Rafiei. 2023. DIN-SQL: Decomposed In-Context Learning of Text-to-SQL with Self-Correction. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023.
- Pourreza and Rafiei (2024) Mohammadreza Pourreza and Davood Rafiei. 2024. DTS-SQL: Decomposed Text-to-SQL with Small Large Language Models. arXiv preprint arXiv:2402.01117 (2024).
- Rane et al. (2024) Nitin Rane, Mallikarjuna Paramesha, Saurabh Choudhary, and Jayesh Rane. 2024. Business Intelligence and Business Analytics With Artificial Intelligence and Machine Learning: Trends, Techniques, and Opportunities. Techniques, and Opportunities (2024).
- Rivera (2024) Kevin Mejía Rivera. 2024. Developing a Media Watcher Through Business Intelligence Tools: An Approach for Emerging Media Industries. In Gaming, Entertainment, and Media Conference. 1–4.
- Robertson and Sparck Jones (1976) SE Robertson and K Sparck Jones. 1976. Relevance Weighting of Search Terms. Journal of the American Society for Information Science 27, 3 (1976), 129–46.
- Sun et al. (2023) Ruoxi Sun, Sercan Ö Arik, Alex Muzio, Lesly Miculicich, Satya Gundabathula, Pengcheng Yin, Hanjun Dai, Hootan Nakhost, Rajarishi Sinha, Zifeng Wang, et al. 2023. SQL-PaLM: Improved Large Language Model Adaptation for Text-to-SQL (extended). arXiv preprint arXiv:2306.00739 (2023).
- Talaei et al. (2024) Shayan Talaei, Mohammadreza Pourreza, Yu-Chen Chang, Azalia Mirhoseini, and Amin Saberi. 2024. CHESS: Contextual Harnessing for Efficient SQL Synthesis. arXiv preprint arXiv:2405.16755 (2024).
- Tamang et al. (2021) Minsang D Tamang, Vinod Kumar Shukla, Shaista Anwar, and Ritu Punhani. 2021. Improving business intelligence through machine learning algorithms. In 2021 2nd International Conference on Intelligent Engineering and Management. 63–68.
- Team (2024) Qwen Team. 2024. Qwen2.5: A Party of Foundation Models. https://qwenlm.github.io/blog/qwen2.5/
- team ([n.d.]) StarRocks team. [n.d.]. StarRocks. https://github.com/StarRocks/starrocks
- Wang et al. (2023) Bing Wang, Changyu Ren, Jian Yang, Xinnian Liang, Jiaqi Bai, Qian-Wen Zhang, Zhao Yan, and Zhoujun Li. 2023. MAC-SQL: A Multi-Agent Collaborative Framework for Text-to-SQL. arXiv preprint arXiv:2312.11242 (2023).
- Wang et al. (2020) Bailin Wang, Richard Shin, Xiaodong Liu, Oleksandr Polozov, and Matthew Richardson. 2020. RAT-SQL: Relation-Aware Schema Encoding and Linking for Text-to-SQL Parsers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 7567–7578.
- Wang et al. (2017) Chenglong Wang, Alvin Cheung, and Rastislav Bodík. 2017. Synthesizing highly expressive SQL queries from input-output examples. In Proceedings of the 38th Conference on Programming Language Design and Implementation. 452–466.
- Wang et al. (2024) Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder, and Furu Wei. 2024. Multilingual e5 text embeddings: A technical report. arXiv preprint arXiv:2402.05672 (2024).
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H Chi, Quoc V Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems. 24824–24837.
- Wieder and Ossimitz (2015) Bernhard Wieder and Maria-Luise Ossimitz. 2015. The impact of Business Intelligence on the quality of decision making–a mediation model. Procedia Computer Science 64 (2015), 1163–1171.
- Yu et al. (2018) Tao Yu, Rui Zhang, Kai Yang, Michihiro Yasunaga, Dongxu Wang, Zifan Li, James Ma, Irene Li, Qingning Yao, Shanelle Roman, Zilin Zhang, and Dragomir R. Radev. 2018. Spider: A Large-Scale Human-Labeled Dataset for Complex and Cross-Domain Semantic Parsing and Text-to-SQL Task. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing. 3911–3921.
- Zeng et al. (2016) Zhong Zeng, Mong-Li Lee, and Tok Wang Ling. 2016. Answering Keyword Queries involving Aggregates and GROUPBY on Relational Databases. In Proceedings of the 19th International Conference on Extending Database Technology. 161–172.

### v2

@@
-SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models
+SiriusBI: A Comprehensive LLM-Powered Solution for Data Analytics in Business Intelligence
 Abstract.
-The rapid advancement of AI technologies, particularly Large Language Models (LLMs), is establishing a new paradigm for Business Intelligence (BI).
-Despite the emergence of pioneering work in enhancing BI systems with LLMs, we have identified the following three issues when deployed in real industrial scenarios: interaction limitations, performance bottlenecks, and functionality deficiencies.
-In this paper, we present SiriusBI, an end-to-end business intelligence system that is designed to address the three issues simultaneously.
-First, we propose an intelligent and application-oriented module called multi-round dialogue with querying, which aims to overcome the prevalent interaction limitations in current BI solutions.
-Next, to mitigate the performance bottlenecks caused by scenario migration, we introduce two SQL generation methods that strike a balance between accuracy and deployment costs.
-Finally, to tackle the practical challenges posed by functionality deficiencies, we develop an end-to-end workflow that covers the entire BI process, ensuring that SiriusBI delivers a robust and complete set of functionalities.
-As an independent cloud service in Tencent’s data platform, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients.
-Experiments on real-world datasets and practical applications in industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI.
-Remarkably, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
+With the proliferation of Large Language Models (LLMs) in Business Intelligence (BI), existing solutions face critical challenges in industrial deployments: functionality deficiencies from legacy systems failing to meet evolving LLM-era user demands, interaction limitations from single-round SQL generation paradigms inadequate for multi-round clarification, and cost for domain adaptation arising from cross-domain methods migration.
+We present SiriusBI, a practical LLM-powered BI system addressing the challenges of industrial deployments through three key innovations: (a) An end-to-end architecture integrating multi-module coordination to overcome functionality gaps in legacy systems; (b) A multi-round dialogue with querying mechanism, consisting of semantic completion, knowledge-guided clarification, and proactive querying processes, to resolve interaction constraints in SQL generation; (c) A data-conditioned SQL generation method selection strategy that supports both an efficient one-step Fine-Tuning approach and a two-step method leveraging Semantic Intermediate Representation for low-cost cross-domain applications.
+Experiments on both real-world datasets and public benchmarks demonstrate the effectiveness of SiriusBI.
+User studies further confirm that SiriusBI enhances both productivity and user experience.
+As an independent service on Tencent’s data platform, SiriusBI is deployed across finance, advertising, and cloud sectors, serving dozens of enterprise clients.
+It achieves over 93% accuracy in SQL generation and reduces data analysts’ query time from minutes to seconds in real-world applications.
 PVLDB Reference Format:
-Jie Jiang,
-Haining Xie,
-Yu Shen,
-Zihan Zhang,
-Meng Lei,
-Yifeng Zheng,
-Yide Fang,
-Chunyou Li,
-Danqing Huang,
-Wentao Zhang,
-Yang Li,
-Xiaofeng Yang,
-Bin Cui,
-Peng Chen
-. PVLDB, 14(1): XXX-XXX, 2020.
+PVLDB, 14(1): XXX-XXX, 2020.
 doi:XX.XX/XXX.XX
@@
 PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
+The source code, data, and/or other artifacts have been made available at https://github.com/TencentBigData-SiriusAI/SiriusBI.
 1. Introduction
-Business Intelligence (BI) (Negash, 2004; Wieder and Ossimitz, 2015) is a crucial application scenario in the data field, comprising a comprehensive suite of methodologies, tools, and infrastructures designed to collect, integrate, analyze, and present an organization’s raw data to generate actionable insights for informed decision-making.
-The primary objective of BI is to generate actionable insights that facilitate informed decision-making.
-BI systems are extensively utilized across various sectors, including finance (Nuseir, 2021), environment (Hegger and Saraceni, 2024), and social media (Chen et al., 2024a; Rivera, 2024), which significantly enhances the decision-making process through the provision of real-time analytics and reporting capabilities.
-A typical BI system comprises several key components: a data management system that stores, processes, and aggregates vast amounts of data; analytics algorithms that transform this data into actionable insights; and visualization tools that present the information in intuitive and user-friendly formats.
-These components collaboratively function to streamline business processes, enhance customer experiences, reduce operational costs, and mitigate risks.
-Recent advancements in LLMs have sparked significant interest in ChatBI (Lian et al., 2024) — an integrated BI system enhanced by conversational AI.
+Business Intelligence (BI) (Negash, 2004; Wieder and Ossimitz, 2015) is a crucial application scenario in the data field, comprising a comprehensive suite of methodologies, tools, and infrastructures designed to collect, integrate, analyze, and present raw data from an organization to generate actionable insights for informed decision-making.
+BI systems are extensively used in various sectors, including finance (Nuseir, 2021), environment (Hegger and Saraceni, 2024), and social media (Chen et al., 2024a; Rivera, 2024), which significantly improves the decision-making process through the provision of real-time analytics and reporting capabilities (Wang et al., 2025; Li, 2024).
+A typical BI system comprises several key components: a data management module that stores, processes, and aggregates vast amounts of data; analytic algorithms that transform the data into actionable insights; and visualization tools that present the information in intuitive and user-friendly formats.
+Among these, data analytics plays a crucial role in providing decision-making support, directly determining the correctness and appropriateness of decisions.
+Recent advancements in LLMs (Lian et al., 2024; Ke et al., 2024; Zhang et al., 2024) have sparked significant interest in ChatBI — a new paradigm supported by natural language interfaces (Affolter et al., 2019).
 Concurrently, the demand for a fully integrated and efficient ChatBI solution is surging, driven by the need of a more intuitive and accessible mode of data interaction.
-This evolution promises to transform how users engage with data, making insights more readily available and actionable.
-To meet the growing demand for big data analysis and decision-making in BI, the data community has proposed numerous effective approaches.
-While NL2SQL is the core task in most BI systems, traditional methods often rely on the schema-based approaches (Hristidis and Papakonstantinou, 2002; Luo et al., 2007; Zeng et al., 2016; Hristidis et al., 2003) or parsing-based approaches (Iyer et al., 2017; Wang et al., 2017; Li and Jagadish, 2014; Popescu et al., 2004) to generate SQL queries from natural language.
-Recently, there has been a growing trend in implementing NL2SQL powered by LLMs, utilizing either prompt engineering (Pourreza and Rafiei, 2023; Wang et al., 2023) or model training techniques (Li et al., 2024; Pourreza and Rafiei, 2024).
-However, these methods typically focus on Single-Round Dialogue (SRD) scenarios, which limits their effectiveness in real-world applications where user interactions are typically more complex.
-To address the limitations of NL2SQL in BI contexts, the NL2BI (Lian et al., 2024) task has been introduced.
-This task aims to generate SQL queries based on Multi-Round Dialogues (MRD), allowing for a comprehensive understanding of user intentions.
-However, when implementing existing work in industrial BI systems and applying it in real-world scenarios, we identified the following three challenges:
-C1: Interaction Limitations.
-ChatBI aims to leverage MRD to assist users in data analysis, thereby lowering the barrier to data usage and enhancing overall data utilization efficiency.
-However, existing public solutions still face challenges in supporting MRD due to real-world interaction modes.
-On one hand, much of the research (Wang et al., 2023; Talaei et al., 2024; Li et al., 2024; Pourreza and Rafiei, 2024, 2023) in the community concentrates on enhancing the performance of the NL2SQL task.
-While this has advanced the field to some extent, the interaction mode remains restricted to SRD, which imposes strict constraints on how users express their needs.
-This limitation ultimately undermines the primary objective of ChatBI systems: to facilitate universally accessible data analysis.
-On the other hand, leading technology teams in industry have introduced ChatBI (Lian et al., 2024) that support MRD and offered feasible methods for implementing multi-round interactions.
-Nevertheless, these solutions also impose limitations on user interaction modes in practical applications, catering only to fixed question-answer formats.
-Furthermore, they tend to be less effective in BI scenarios that require deep domain expertise.
-C2: Performance Bottlenecks.
-The NL2SQL task is crucial for BI because it directly impacts the accuracy of data retrieval during analysis.
-In practical industrial scenarios, existing NL2SQL methods often exhibit suboptimal performance when applied across diverse contexts.
-For example, a LLM fine-tuned on data specifically annotated for the financial domain performs poorly when applied to the advertising domain.
-The process of preparing and standardizing data for each specialized field, followed by fine-tuning, incurs significant costs, which are often impractical in industrial applications.
-C3: Functionality Deficiencies.
-Traditional BI systems are typically comprehensive solutions for data analysis, encompassing multiple interrelated modules such as data storage, SQL generation, and data insights.
-However, existing LLM-based BI systems fall short in providing essential functionalities, including efficient intent understanding, adaptive domain-specific knowledge, and intelligent data insights.
-For example, Mac-SQL (Wang et al., 2023) and CHESS (Talaei et al., 2024) primarily focus on NL2SQL tasks, while Lian et al.
-(2024) propose using the Apache SuperSet(Apa, [n.d.]) tool directly after NL2SQL, neglecting the components of domain knowledge management module and deep data insight.
-Unlike traditional BI systems, users expect LLMs to enhance the entire ChatBI workflow, which drives a technological paradigm shift that improves both application efficiency and user experience.
-A purely NL2SQL-based approach or scenario-constrained MRD functionality is inadequate to meet the real-world needs of BI applications.
-This gap presents an opportunity to develop a more comprehensive and intelligent ChatBI system.
-To address the aforementioned challenges, we propose SiriusBI, which implements an end-to-end solution for BI scenarios.
-This system leverages the capabilities of LLMs to empower various submodules, thereby enhancing both efficiency and user experience in data analysis.
-Specifically, for the issue of interaction limitations (C1), we introduce the MRD-Q (Multi-Round Dialogue with Querying) module.
-Unlike the MRD functionality proposed by Lian et al.
-(2024), MRD-Q incorporates an intent clarification module that disambiguates user queries through follow-up questions.
-This approach enables the system to accurately identify the user’s true intent, even when the initial query is flawed or ambiguous, thereby facilitating precise responses, as shown in Figure 1.
-Furthermore, we employ a Retrieval-Augmented Generation (RAG) framework that leverages both business-domain knowledge and metadata, ensuring our system’s adaptability across various domains.
-Consequently, this reduces barriers for users and significantly enhances the overall user experience.
-For the issue of performance bottlenecks (C2), we propose two implementation strategies in SQL generation.
-First, for domain-specific NL2SQL tasks, we introduce an efficient and automated process for constructing training data to improve the performance of LLMs.
-Second, for scenarios that require transferability, we present a two-step solution that integrates a Semantic Model with a SQL Model.
-This approach leverages domain-specific knowledge and metadata, ensuring high accuracy in SQL generation without the need of additional training.
-Furthermore, it reduces the costs associated with transferring between different scenarios and enhances the generalization of NL2SQL across various domains.
-For the issue of functionality deficiencies (C3), SiriusBI provides a comprehensive end-to-end workflow.
-Unlike existing works that mainly focus on the NL2SQL task and the expansion of MRD modules, SiriusBI introduces two essential modules specifically designed for ChatBI : the Knowledge Management module and the Data Insights module.
-The knowledge management module is responsible for storing, processing, and utilizing domain-specific expertise and data to support the specialized knowledge requirements of various submodules.
-Meanwhile, the Data Insights module transcends traditional data visualization and report generation by providing critical key functionalities such as automatic data preparation, complex task planning, and the application of insight tools.
-These capabilities provide users with more comprehensive and in-depth support for their data analysis needs.
+This evolution promises to transform how users engage with data, making insights more available and actionable.
+To meet the growing demand for big data analytics and decision-making in BI, the data community has proposed numerous effective approaches.
+However, when applying existing work in real-world BI scenarios, we identify the following three challenges:
+C1: Functionality Deficiencies.
+While traditional business intelligence systems (Chaudhuri et al., 2011) integrate core components spanning data management, SQL generation, and insight discovery to form complete analytics pipelines, their reliance on heuristic rules and conventional AI/ML techniques limits generalization ability in dynamic scenarios.
+Although LLM-based methods have advanced task-specific performance, few offer comprehensive BI capabilities comparable to their traditional counterparts.
+For example, MAC-SQL (Wang et al., 2023b) and CHESS (Talaei et al., 2024) optimize NL2SQL accuracy but treat SQL execution as terminal outputs, neglecting downstream tasks like attribution analysis.
+While Lian et al.
+(Lian et al., 2024) extend their pipeline with Apache Superset for visualization, they fail to introduce knowledge bases to support dynamic grounding of domain-specific context, a critical requirement for real-world BI adaptation (Cheng et al., 2009).
+This functional fragmentation forces users to manually coordinate tools (e.g., SQL editors, dashboard platforms, knowledge retriever), which imposes significant cognitive load and reduces operational efficiency, as evidenced by industry report (Kurt Schlegel, 2023).
+C2: Interaction Limitations.
+In the context of ChatBI, the NL2SQL task is becoming increasingly vital, as it facilitates seamless interaction between natural language queries and structured data retrieval, thereby enhancing the efficiency and accuracy of data analytics.
+The evolution of NL2SQL techniques reveals a critical architectural mismatch: while traditional methods (schema-based (Hristidis and Papakonstantinou, 2002; Luo et al., 2007; Zeng et al., 2016; Hristidis et al., 2003) or parsing-based (Iyer et al., 2017; Wang et al., 2017; Li and Jagadish, 2014; Popescu et al., 2004)) and modern LLM-driven approaches (prompt engineering (Pourreza and Rafiei, 2023; Wang et al., 2023b) or fine-tuning techniques (Li et al., 2024c; Pourreza and Rafiei, 2024)) predominantly optimize for Single-Round Dialogue (SRD) precision.
+This SRD-centric paradigm introduces a significant continuity gap in Multi-Round Dialogues (MRD): real-world BI workflows often require iterative investigation through successive queries, where later queries tend to omit previously provided contextual information, resulting in semantic ambiguity beyond the initial query.
+For instance, in the MRD NL2SQL task illustrated in Figure 1, the user issues three queries; notably, the second and third queries omit the time condition “2023” because it was specified in the first query.
+Single-round NL2SQL approaches demand nearly perfect input specificity, which explains its failure to generate correct SQL statements for the second and third queries in Figure 1.
+Worse still, due to the intricate nature (Liu et al., 2024) of MRD, few approach has been devoted to addressing this task.
+Lian et al.’s MRD solution (Lian et al., 2024) is the first attempt towards this task.
+Nevertheless, their solution is absent of user-guided clarification loops for intent resolution and domain-grounded dialogue act modeling.
+As demonstrated in Figure 1, the basic MRD approach still exhibit performance degradation beyond the first dialogue round.
+C3: Cost for Domain Adaptation.
+Cross-domain deployment of NL2SQL models faces the challenge of cost surges in domain knowledge transfer, primarily caused by insufficient model generalization capability.
+Structural differences in database schema across domains (e.g., nested tables in finance vs.
+wide tables in advertising) necessitate repetitive model adaptation (Liu et al., 2024), while semantic gaps between industry-specific operators (e.g., financial window functions vs.
+e-commerce promotional rules) exacerbate logical deviations in SQL generation (Li et al., 2024a).
+Critically, domain knowledge transfer relies heavily on expert-annotated data, with manual annotation costs growing with domain complexity (Li et al., 2024b).
+Our real-world deployment statistics show that direct model migration leads to business logic errors in approximately two-thirds of generated SQL queries.
+Meanwhile, adapting models through traditional fine-tuning requires 5.5 person-days on average to label 200 seed queries within existing databases—forming critical bottlenecks for enterprise-level scalability.
+To address the aforementioned challenges, we propose SiriusBI, which implements a comprehensive LLM-powered solution for ChatBI scenarios.
+This system leverages the capabilities of LLMs to empower various modules, thereby enhancing both the efficiency and user experience in data analytics.
+Specifically, for the issue of functionality deficiencies (C1), SiriusBI introduces an end-to-end integrated architecture that seamlessly orchestrate core modules including knowledge management, multi-round dialogue analysis, SQL generation, and data insight provision, thereby ensuring a closed-loop pipeline from natural language queries to final decision-making reports.
+For the issue of interaction limitations (C2), we introduce the MRD-Q (Multi-Round Dialogue with Querying) module.
+As a supplement to the basic multi-round dialogue analysis module proposed by Lian et al.
+(Lian et al., 2024), MRD-Q incorporates an intent querying module to clarify user queries through follow-up questions.
+This approach enables the system to accurately identify the user’s true intent, even when the initial query is incomplete or ambiguous, thus facilitating precise responses, as shown in Figure 1.
+To enable economic domain adaptation (C3), SiriusBI introduces a strategy switching mechanism that dynamically selects between one-step and two-step SQL generation paradigms based on data conditions.
+This mechanism optimizes the trade-off between computational cost and performance, ensuring efficient adaptation to diverse business domains.
+To address structural and semantic disparities across domains, the one-step method employs a supervised fine-tuning (SFT) method enhanced by an automated data generation process pipeline.
+This pipeline minimizes manual annotation costs by leveraging SQL logs and LLM-based reverse engineering to synthesize high-quality training data, greatly reducing reliance on labor-intensive expert labeling.
+For cold-start scenarios, the two-step method eliminates SFT by combining domain knowledge retrieval with semantic intermediate representation (SIR), achieving robust performance across diverse industries without domain-specific tuning.
+The generalization ability of our system is further enhanced by a knowledge base with hybrid storage strategies, specially designed for BI application scenario.
 Contributions. The main contributions of this paper can be summarized as follows:
-We propose SiriusBI, an end-to-end LLM-based ChatBI system designed to provide better data analysis services and lower the usage threshold for users.
-To the best of our knowledge, SiriusBI is the first comprehensive end-to-end ChatBI system that simultaneously addresses the aforementioned three challenges.
-So far, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients.
-To extend the interaction modes of ChatBI systems, we introduce MRD-Q, a dialogue system designed to ask clarifying questions when user queries are unclear or ambiguous.
-This improves the accuracy of intent understanding, lowers the barrier to user entry, and ultimately enhances the overall user experience.
-To enhance the transferability of ChatBI systems across different industrial scenarios, we propose two implementation approaches for SQL generation: single-step generation and two-step generation.
-The two-step generation method significantly improves the transferability of ChatBI and provides a new paradigm for SQL generation in the industry.
-Extensive experiments on real-world datasets and industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI.
-Notably, when applied in industry, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
-2. Preliminary
-In this section, we introduce some key concepts including BI, ChatBI, NL2SQL, knowledge management and data insight.
-2.1. BI and ChatBI
-BI refers to the technologies, applications, and practices involved in the collection, integration, analysis, and presentation of business information.
-The primary goal of BI is to improve decision-making with business organizations.
-Traditional BI systems typically rely on data warehousing, data mining, and reporting tools, which enable organizations to transform raw data into meaningful insights.
-ChatBI represent an evolution of traditional BI systems by integrating natural language processing (NLP) capabilities.
-This integration allows users to interact with BI tools using conversational interfaces.
-ChatBI systems aim to democratize data access by enabling non-technical users to query and analyze data through natural language queries.
-These systems leverage advanced NLP techniques to accurately understand user intent and generate relevant responses, thereby enhancing the overall user experience and allow organizations to make data-driven decisions more accessible and actionable.
-2.2. NL2SQL Task
-NL2SQL is a task in the field of natural language processing and database management that focuses on converting natural language queries into SQL queries.
-In the specific task of SQL generation, we first receive the user’s query intent .
-Then, based on the relevant metrics and dimensions within the user’s intent, and in conjunction with the database schema information , and the SQL query result can be generated as,
-where is the function that generates the SQL query result .
-The accuracy and efficiency of NL2SQL systems are critical for the effectiveness of ChatBI systems, as they directly impact the quality of the generated insights.
+Practical ChatBI System: SiriusBI integrates four core functionalities — knowledge management, multi-round dialogue with querying, SQL generation, and data insight — into a unified system.
+It has been deployed across Tencent’s finance, advertising, and cloud businesses departments, serving dozens of enterprise clients, reducing data analysis latency from minutes to seconds.
+New Dialogue Analysis Mechanism: We propose a new dialogue analysis mechanism named MRD-Q, consisting of semantic completion, knowledge-guided clarification, and proactive querying processes.
+It effectively addresses issues raised by ambiguous or incomplete queries in real-world scenarios.
+Economic Domain Adaptation Strategy: We propose a dynamic strategy selection mechanism for SQL generation conditioned on data characteristics.
+To reduce annotation costs for fine-tuning, we set up an automated data synthesis pipeline.
+To enhance cross-domain SQL syntax accuracy and business logic consistency, we use Semantic Intermediate Representation (SIR) in prompt engineering.
+Multi-Round NL2SQL Benchmark: We construct MRD-BIRD —a dataset containing 96 multi-round dialogue sequences (3–5 interaction rounds)— and open-source the data to advance NL2SQL research in multi-round dialogue contexts.
+Proven Effectiveness: Extensive experiments on both academic benchmarks and real-world datasets demonstrate the effectiveness of SiriusBI.
+Notably, when applied in industry, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 93% for Tencent Advertisement, and 96% for Tencent Cloud.
+User studies further confirm that SiriusBI enhances both productivity ad user experience.
+2. Related Work
+2.1. Business Intelligence
+The integration of AI and Machine Learning (ML) has transformed data-driven decision-making in BI (Rane et al., 2024).
+ML and Predictive Analytics enhance forecasting and pattern recognition in large datasets (Lee et al., 2022; Tamang et al., 2021).
+NLP and Text Analytics extract insights from unstructured data, improving customer service and decision-making (Kaur et al., 2019; Ittoo et al., 2016).
+The rise of the ChatBI paradigm, which has significantly advanced business intelligence and taken a major step towards data democratization (Linkon et al., 2024), is largely driven by the development of advanced AI models such as BERT (Devlin et al., 2019) and ChatGPT (team, 2022), which enable intelligent data interpretation and improved user interactions.
+Among related works, BI-REC (Meduri et al., 2021) is a pioneering system that recommends BI applications to help users achieve their data analysis goals.
+Similar to our method, BI-REC includes a step for predicting BI intents.
+However, BI-REC focuses on leveraging historical state information to predict possible future BI intents, whereas our approach aims to assess whether the current user intent is sufficiently clear and precise to generate an accurate SQL query.
+Lian et al.
+(Lian et al., 2024) are the first to recognize the importance of the NL2BI task in practical ChatBI applications and propose an efficient method to address it.
+We draw inspiration from their dialogue analysis technique to obtain semantically complete queries.
+The key differences between our approach and theirs are: (1) we leverage LLMs instead of smaller pre-trained models to analyze query completeness; (2) we incorporate domain-specific knowledge to help the LLM better understand user intentions; and (3) we enable active clarification questions directed at users when the LLM determines that the query requires further disambiguation.
+Lian et al.’s method does not consider the latter two aspects.
+2.2. NL2SQL Methods
+NL2SQL is a core task in ChatBI systems and can be broadly categorized into prompt-based and training-based methods (Zhou et al., 2024).
+Recently, Large Language Models (LLMs) like Codex (team, 2023a) and Claude 3 (Team, 2024) have shown great promise in NL2SQL, leveraging techniques such as prompt engineering (Giray, 2023), in-context learning (Dong et al., 2022), and chain-of-thought prompting (Wei et al., 2022).
+Prompt-based approaches not only offer new technical paradigms but also reduce costs in industrial settings.
+Notable examples include DIN-SQL (Pourreza and Rafiei, 2023), C3 (Dong et al., 2023), MAC-SQL (Wang et al., 2023b), and SQL-PaLM (Sun et al., 2023), which improve NL2SQL performance on models like ChatGPT and PaLM (Narang and Chowdhery, 2022).
+Before LLMs, NL2SQL solutions mainly relied on training encoder-decoder sequence-to-sequence models (Cai et al., 2021; Cao et al., 2021; Choi et al., 2021; Gan et al., 2021; Wang et al., 2020a), requiring large amounts of database and SQL data.
+The advent of LLMs has introduced new training paradigms; for instance, Codes (Li et al., 2024c) proposes a training method tailored for NL2SQL that boosts LLM performance.
+However, some studies (Maamari et al., 2024; Pourreza et al., 2024; Caferoğlu and Ulusoy, 2024) fine-tune LLMs like ChatGPT (team, 2022) and Gemini (Team et al., 2023) to further improve results, though this often incurs high practical costs.
+3. Method
+3.1. Overview
+The overall framework of SiriusBI is illustrated in Figure 2 and consists of four core modules:
+(1) The Knowledge Management module consists of four types of data – database metadata, domain-specific knowledge, task-related materials (e.g., demonstrations used for in-context learning in LLMs), and historical conversations.
+These data types are stored in a hybrid architecture, ensuring efficient and scalable access to knowledge, supporting the seamless functionality of other modules.
+(2) The Multi-Round Dialogue with Querying (MRD-Q) module is composed of two components: Dialogue Analysis and Intention Querying, which enable multi-turn user-oriented dialogue and knowledge-guided clarification to resolve semantic ambiguity in Conversational BI.
+Unlike existing SRD-centric approaches (Wang et al., 2023b; Pourreza and Rafiei, 2023), MRD-Q introduces follow-up question generation based on domain-specific knowledge, enabling context-aware SQL calibration across 3-5 dialogue turns.
+(3) The SQL Generation module dynamically selects between one-step (fine-tuning based) and two-step (semantic intermediate representation based) processing paradigms according to business scenario data characteristics.
+This dual-mode approach efficiently transforms user intent into SQL queries with optimal cost-effectiveness.
+The one-step method streamlines the process through an automated data preparation pipeline, while the two-step method employs a knowledge-base-driven query rewriting strategy that eliminates model training requirements.
+This architectural design balances operational efficiency with implementation flexibility across diverse application contexts.
+(4) To close the loop from SQL execution to decision support, SiriusBI integrates a Data Insight module that interprets query results through a multi-agent workflow based on the FunctionCall (Kim et al., 2024) and ReAct (Yao et al., 2023) mechanism as shown in Figure 3.
+Specifically, the Planner Agent analyzes user queries, SQL results, and domain knowledge to decompose tasks into subtasks and issue instructions; the Data Preparation Agent identifies missing data requirements and generates SQL queries to retrieve necessary information; the Tool Execution Agent selects specialized tools (Forecast/Diagnosis/Attribution Tools, implemented via reproduced existing works (Box et al., 2015; Hundman et al., 2018; Bhagwan et al., 2014; Li et al., 2019; Lundberg and Lee, 2017)) to process data, with final outputs consolidated by the Planner Agent for response generation.
 3.2. Knowledge Management
-Knowledge Management (KM) involves systematic processes within an organization aimed at optimizing the use of knowledge.
-In most BI and ChatBI systems, KM is structured into three modules: knowledge processing, knowledge storage, and knowledge retrieval.
-The critical formula for knowledge retrieval is:
-where represents the structured knowledge from retrieval, represents the user’s query, denotes the stored knowledge, and is the function that facilitates the retrieval of knowledge from storage.
-2.4. Data Insight
-Data insights is a fundamental component of BI and ChatBI systems, designed to transform raw data into actionable insights through a series of structured processes: data aggregation, data analysis, insight generation, and visualization.
-Data Aggregation: Combine data from multiple sources into a comprehensive dataset.
-Data Analysis: Apply statistical and machine learning techniques to identify patterns and trends within the aggregated data.
-Insight Generation: Translate the results of data analysis into meaningful insights that can inform business decisions.
-Visualization: Present insights through intuitive visual formats to enhance understanding and facilitate decision-making.
-3. System design
-In this section, we introduce the overall architecture of SiriusBI.
-Compared with the current ChatBI systems (Lian et al., 2024), SiriusBI provides a more comprehensive workflow, especially in the knowledge management and data insight.
-3.1. Overview
-The overall framework of SiriusBI is illustrated in Figure 2 and consists of five key modules: (1) The Knowledge Management subsystem which stores, processes, and retrieves pertinent database knowledge, domain-specific expertise, and demonstrations (i.e., examples used for context learning in LLMs) that support the functionality of the other modules.
-(2) The Dialogue Analysis module, which accepts user queries and handles the interaction through multiple-round dialogues.
-(3) The Intention Querying module, which clarifies the user’s intent by generating follow-up questions based on the dialogue history.
-(4) The SQL Generation module, which converts the fully understood user intent into SQL queries.
-(5) The Data Insight module, which automatically prepares data, plans tasks, attributes issues, and predicts trends based on the data retrieved through NL2SQL queries.
-3.2. Workflow
-The workflow of SiriusBI is detailed as follows: When a user inputs a question in a BI scenario, for example, ”Company A’s year-to-date consumption compared year-over-year?” (), in the first stage, the Dialogue Analysis module evaluates whether the semantic content of the user’s input is complete, i.e., whether it includes both ”dimensions” and ”metrics”.
-If the input is determined to be semantically complete, it proceeds to the second stage, the Intention Querying module.
-However, if the user’s input is missing either dimensions or metrics, the module uses the conversation history to complete the semantics of the current question before proceeding to the next stage.
-For instance, if the user subsequently asks, ”How about Company B?” (), the Dialogue Analysis module will complete the question to ”Company B’s year-to-date expenditure compared year-over-year?” ().
-In the second stage, the Intention Querying module recalls relevant information from the knowledge base based on the input question.
-Notably, to incorporate domain knowledge, improve system generalization, and enhance data insights, we have developed a specialized Knowledge Management module, which is elaborated in 4.1.
-After knowledge retrieval, we recall relevant fields from the database.
-If there is still ambiguity or confusion among the recalled fields, the Intention Querying module prompts the user for clarification to ensure accurate input for the SQL generation stage, thereby minimizing performance loss due to error propagation in the pipeline.
-In the third stage, the SQL generation module receives the semantically completed user query from the previous stage and selects the appropriate tables for each query.
-Typically, this involves matching the top-k tables using database table information.
-To improve the adaptability of SiriusBI to various scenarios and requirements, we propose two optional generation methods: 1) An one-step generation method, where we construct specific datasets via an automated process for and perform Supervised Fine-Tuning (SFT) on tasks that require strong domain expertise but lower generalization; 2) A two-step SQL generation method, where we structure user queries via knowledge retrieval and then translate inputs into SQL using a fixed LLM.
-This method ensures fast deployment of SQL generation on unseen scenarios.
-By parallel use of these two generation methods, we maximize the accuracy of the generated SQL queries, ensuring reliable and relevant results for users.
-In the fourth stage, the Data Insight module receives the SQL queries and their corresponding results from the third stage.
-Based on the user’s additional insight requirements, the module performs automatic data preparation, complex task planning, insight tool execution, and finally generates corresponding reports.
-In the following section, we provide more in-depth exploration of the design and functionalities of each module.
-4. System description
-In this section, we explore various components of SiriusBI in detail.
-We begin with the Knowledge Management module in Section 4.1, which serves as a critical foundation that provides essential knowledge support for other parts of the workflow.
-Next, Section 4.2 focuses on Multi-round Dialogue with Querying, which consists of both the Dialogue Analysis and Intention querying module.
-In Section 4.3, we discuss SQL Generation, highlighting both the one-step and two-step SQL generation schemes.
-Finally, we introduce the Data Insight module in Section 4.4.
-4.1. Knowledge Management
-In industrial scenarios, the application of Business Intelligence (BI) often requires a substantial amount of domain-specific knowledge, including table metadata, business knowledge, and LLMs chat history, etc.
-To address the complex knowledge requirements, SiriusBI establishes the Knowledge Management module by constructing a multidimensional knowledge graph.
-This knowledge graph comprises several key dimensions: 1) metadata information, which includes table schema, column values; 2) business knowledge, which includes domain-specific fields, term definitions, and relevant business context; 3) demonstration information utilized for prompting; 4) conversation data.
-Together, these dimensions establish a robust foundation for the knowledge management module.
-The module also supports DSL (Domain-Specific Language) data generation, which includes functionalities such as query rewriting and DSL prompting.
-Furthermore, it facilitates API data retrieval and provides essential metadata and business knowledge, thereby supporting comprehensive data insights.
-This versatility makes the knowledge management module a vital component in various data-driven applications.
+The application of BI often requires substantial amount of domain-specific knowledge.
+To address the knowledge requirements, SiriusBI constructs and maintains a comprehensive knowledge base within the knowledge management module.
+This knowledge base is designed to support efficient data retrieval, enhance decision-making processes, and facilitate seamless integration with BI tools.
+The knowledge base primarily constitutes four types of data: (1) Database Metadata: This includes essential information such as table names, column types, column names, and other details of the database schema.
+(2) Domain-Specific Knowledge: This encompasses explanations of certain fields, term definitions, and relevant business context.
+It provides the necessary background information to interpret and utilize the data effectively within the specific domain.
+(3) Task-Specific Materials: These are groups of prompts, demonstrations and other resources designed for different tasks.
+They are tailored to guide users in performing specific operations or analyses within the BI system.
+(4) Historical Conversations: This includes logs of past interactions and queries, which can be used to complete semantic for multi-round dialogue analysis.
 3.2.1. Knowledge Storage
-For knowledge storage, we use EasyGraph (Gao et al., 2023) to construct a multidimensional knowledge graph.
-Figure 3 shows the graph stucture for storing knowledge related to data tables.
-Each node is a data structure that contains node types, descriptions, embeddings, and other information.
-The node types in this graph include databases, tables, columns, (column) values, UDFs (complex metric calculation functions such as month-over-month), terms (explanations of professional terms), and aliases (alternative names for other nodes).
-For columns, the node has specific attributes like whether it is a dimension or measure.
-The relationships between these nodes are represented as edges, forming a comprehensive knowledge storage structure used in SiriusBI.
-By defining graph structures and enhancing query efficiency through advanced graph algorithms, we not only organize knowledge effectively but also facilitate the rapid retrieval of relevant information during queries.
-4.1.2. Knowledge Processing
-Before storing the knowledge, it’s essential to process various types of data to ensure effective storage and management.
-We provide standard guidelines for the business teams to follow.
-To create a new node, domain experts first manually extract the node attributes based on specific business scenarios.
-Then, for knowledge with missing attributes, we complete it through metadata enhancement (Choudhury et al., 2023).
-This enhancement occurs on two levels: 1) automatic context completion using AI techniques, and 2) backfilling missing metadata through data engineering methods.
-Together, these processes ensure that the knowledge graph is ready for effective utilization.
-4.1.3. Knowledge Retrieval
-The knowledge retrieval process is divided into two main phases: coarse retrieval and fine retrieval. Figure 4 shows a brief overview of the retrieval process.
-Coarse Retrieval.
-Initially, we use Elasticsearch (Elasticsearch, 2018) for coarse retrieval without segmenting the original query.
-Specifically, we deploy a hybrid search on both the lexical and semantic levels (Wang et al., 2024).
-To achieve a high recall rate, we allow for a larger recall set, temporarily prioritizing recall over precision.
-For each recalled node, if it is an alias or a value, we trace back to at least the column, term, or UDF level.
-For example, for SQL generation, this process yields all potentially relevant column names for fine retrieval.
-Fine Retrieval.
-Based on the results from coarse retrieval, we perform a secondary search on the sub-graph obtained from the coarse retrieval by selecting the top-k nodes ranked by a reranking model (Chen et al., 2024b), which significantly reduces the number of candidate nodes.
-In addition to reranking, for SQL generation, we apply a LLM to filter out columns if it is not relevant to the user query.
+We observe that certain types of knowledge entries exhibit dependencies, while others remain relatively independent.
+In our application scenario, the relationships between database metadata and relevant domain-specific knowledge can be represented as a directed acyclic graph as illustrated in Figure 4.
+To ensure comprehensive retrieval of all related knowledge, it is essential to explicitly capture and represent these relationships.
+While graph databases (Gao et al., 2023) offer a natural solution for storing such relational knowledge, we opt against their use due to their inherent complexity and overhead, which exceed the requirements of our system.
+Instead, we adopt a lightweight approach by augmenting knowledge entries with two dedicated fields: Anc.
+Label, which specifies the label of the ancestor entry, and Anc.
+Name, which encodes the concrete value of the ancestor entry.
+This design, as depicted in Figure 5 (right table), preserves the relational semantics necessary for knowledge retrieval while avoiding the operational and computational complexities associated with native graph storage.
+This approach strikes a balance between functionality and efficiency, aligning with the practical needs of our system.
+For independent knowledge entries, we simply store a description of the entry and, if applicable, a demonstration or example.
+This straightforward approach is sufficient for entries that do not have dependencies or relationships with other knowledge items, as shown in Figure 5 (left table).
+3.2.2. Knowledge Extraction
+The objective of the knowledge extraction process in SiriusBI is to efficiently and accurately extract relevant information from the knowledge base.
+We employ different knowledge extraction strategies based on the storage strategy described earlier.
+All queries must undergo at least one round of knowledge retrieval, which is divided into two main phases: coarse retrieval and fine retrieval.
+This two-phase approach ensures a balance between recall and precision, enabling the system to first identify a broad set of potentially relevant knowledge and then refine the results to meet the specific needs of the task at hand.
+In the coarse retrieval phase, the primary goal is to maximize recall, ensuring that a comprehensive set of potentially relevant knowledge entries is retrieved.
+This is particularly important for tasks such as SQL generation, where the system needs to identify all possible column names or metadata that might be relevant to the user’s query.
+To achieve this, we employ a mixed-metric search strategy that operates on both lexical and semantic levels (Wang et al., 2024).
+Lexical search captures exact matches or close variants of the query terms, while semantic search leverages embeddings to identify conceptually related knowledge entries, even if they do not share exact lexical overlap.
+For this phase, we utilize Elasticsearch (Elasticsearch, 2018), a widely adopted search engine known for its scalability and efficiency.
+The query is processed without segmentation, allowing for a broader initial recall set.
+For knowledge entries that exhibit hierarchical relationships (e.g., those with Anc.
+Name fields), we implement a multi-round retrieval strategy inspired by multi-hop graph traversal.
+Specifically, for each entry retrieved in the initial search, if its Anc.
+Name field is non-empty, we recursively retrieve the ancestor entries.
+This process continues iteratively until no further ancestor entries are found.
+This approach ensures that the retrieval process captures not only the directly relevant knowledge but also the contextual and hierarchical relationships that are critical for tasks such as SQL generation or business rule interpretation.
+To illustrate this process, consider the right table in Figure 4.
+The retrieval begins with the entry “Campus Hiring”, whose Anc.
+Name field contains the ancestor entry “Campus Recruitment”.
+The system then retrieves “Campus Recruitment”, whose Anc.
+Name field includes “Recruitment Channels”.
+Next, “Recruitment Channels” is retrieved, which further links to “Personnel Table”.
+Finally, “Personnel Table” is retrieved, whose Anc.
+Name field points to “Company HR Database”.
+Since “Company HR Database” is not present in the knowledge base, the retrieval process terminates.
+This example demonstrates how the multi-hop traversal strategy effectively captures a chain of related knowledge entries, ensuring comprehensive retrieval of contextually relevant information.
+The coarse retrieval phase typically yields a large set of candidate knowledge entries.
+To refine these results, we perform a fine retrieval phase, which focuses on improving precision by re-ranking and filtering the candidate set.
+We employ a state-of-the-art re-ranking model, BGE (Chen et al., 2024b), to rank the candidate nodes based on their relevance to the query.
+For tasks such as SQL generation, we further refine the results by leveraging a LLM to filter out irrelevant columns.
+The LLM evaluates the semantic relevance of each column to the user query and removes those that do not contribute to the intended SQL query.
+This step ensures that the final set of retrieved knowledge is both concise and highly relevant to the task.
+By combining coarse retrieval, relationship-aware multi-hop traversal, re-ranking, and LLM-based filtering, SiriusBI achieves a robust and efficient knowledge retrieval process.
+This approach not only ensures high recall and precision but also adapts to the complex dependencies and relationships inherent in domain-specific knowledge, making it well-suited for a wide range of BI applications.
 3.3. Multi-Round Dialogue with Querying
-Multi-Round Dialogue with Querying (MRD-Q) is a critical component in SiriusBI that receives and processes user input.
-It primarily supports multi-round dialogue (MRD), intent recognition, and clarification aimed at the user.
-This reduces the usage barrier, enhances the user experience, and minimizes the impact of errors caused by incorrect intent recognition.
-MRD-Q consists of two modules: dialogue analysis and intention querying.
-A detailed introduction is as follows:
+Multi-Round Dialogue with Querying (MRD-Q) is a critical component in SiriusBI that receives and process user inputs for data integrity.
+It is designed to support MRD through automatic semantic completion and user-involved intent querying.
+A detailed introduction is presented in the following sections.
 3.3.1. Dialogue Analysis
-In the industrial sector, BI has a wide range of applications aimed at enhancing user experience in data retrieval, data analysis, and data science.
-In this context, the multi-round dialogue feature in ChatBI plays a crucial role.
-However, most current NL2SQL research (Pourreza and Rafiei, 2023; Wang et al., 2023; Sun et al., 2023; Caferoğlu and Ulusoy, 2024; Li et al., 2024) is designed based on single-round dialogues, which creates a significant gap between this approach and real-world application scenarios.
-Firstly, single-round NL2SQL tasks require highly precise user input, meaning that even minor errors in the input may prevent the generation of correct SQL statements for subsequent data processing tasks.
-This high requirement not only severely diminishes user experience but also raises the barrier to entry for users who may not be familiar with SQL syntax.
-Secondly, data tasks, including data retrieval, are not isolated processes.
-Real-world data analysis or data science tasks often involve complex and detailed work, which inevitably requires extensive data retrieval efforts and even more sophisticated BI functionalities.
-The single-round dialogue approach for data retrieval is evidently inadequate to meet the demands of these real-world application scenarios.
-The multi-round dialogue feature in SiriusBI is built upon the basic principles of dialogue analysis implementation in ChatBI(Lian et al., 2024).
-It determines the input for the SQL Generation module by assessing whether the user’s current input or historical conversation records contain indicators (columns) and dimensions.
-Specifically, the system first determines whether the user’s current input contains the two key pieces of information: indicators and dimensions.
-If both are found, the current input is considered semantically complete, and the workflow proceeds to the next module.
-If either is missing, the system supplements the current input’s semantics by selecting the most recent historical input that includes indicators and dimensions, and then moves to the next module.
-Notably, unlike ChatBI (Lian et al., 2024), SiriusBI does not use traditional small NLP model fine-tuning to determine whether a statement contains indicators and dimensions or to select relevent historical conversations.
-Instead, we leverage the robust semantic understanding capabilities of LLMs along with the domain-specific knowledge in our knowledge base.
-This approach not only facilitates the system’s transition between different vertical domains but also enhances SiriusBI’s adaptability across various fields, which leads to a reduction on time and data construction costs associated with model training.
+The multi-round dialogue analysis mechanism in SiriusBI is built upon the basic dialogue analysis module in previous work(Lian et al., 2024).
+It first determines the semantic integrity of the user query by assessing whether the current input or historical conversations contain metrics and dimensions.
+In the context of business intelligence, dimensions refer to attributes that provide granularity level information of the user query, such as time periods or geographical locations.
+For example, “last quarter” is a time dimension and “in New York” is of location dimension.
+Metrics, on the other hand, are quantitative measurements, such as the number of sales, which might be represented by a column like Total_Sales.
+When a query includes both dimensions and metrics, such as “the number of sales in New York last month”, it is possible to construct a corresponding SQL query.
+Conversely, if a query lacks either a dimension or a metric, such as “How about Los Angeles?”, it becomes challenging to generate a meaningful SQL query.
+Although this approach is somewhat intuitive, it effectively addresses most dialogue patterns encountered in BI scenarios.
+If both are found, the current input is considered of semantic integrity, and the workflow proceeds to the next module.
+If either is missing, the system supplements the current input’s semantics by selecting the most recent historical input that includes metrics and dimensions, and then moves to the next step.
+Unlike previous work, SiriusBI does not fine-tune a PLM (Pre-trained Language Model, e.g., (Lian et al., 2024) adopts ERNIE (Sun et al., 2019)) to determine the semantic integrity or to select relevant historical conversations.
+PLMs, such as BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019), and ERNIE (Sun et al., 2019), are large-scale neural networks pre-trained on vast amounts of text data to capture general language representations.
+While they have shown remarkable performance in various natural language processing tasks, fine-tuning them for specific applications often demands substantial labeled data and computational resources.
+Additionally, fine-tuned models may not generalize well across different domains, limiting their transferability and effectiveness in diverse BI scenarios.
+Instead, we leverage the robust reasoning capability of LLMs to figure out the existence of “metric” and “dimension” and the historical conversations.
 3.3.2. Intention Querying
-To address the interaction limitations, we have added an intent clarification feature to the basic dialogue analysis module, resulting in a comprehensive MRD-Q component.
-The intention querying module is a critical supplement to dialogue analysis and includes the following three steps:
+For multi-round dialogues in real-world scenarios, relying solely on the dialogue history is often insufficient to complete the missing information due to the inherent complexity.
+To address these challenges, we enhance the basic dialogue analysis module by introducing an intent querying feature, resulting in a robust MRD-Q component.
+The intention querying process consists of the following steps:
 Intent Classification.
-When processing information from dialogue analysis, the first step is to reclassify the user’s true intentions.
-In multi-turn dialogues in real-world scenarios, certain metrics or dimensions may be missing, and relying solely on the dialogue history cannot complete the information.
-At this point, the results of intent classification are needed to decide whether to ask the user follow-up questions.
-We use the same LLM for semantic understanding as in Section 4.2.1 and categorize queries into three types.
+We use the same LLM for dialogue analysis in Section 3.3.1 and further categorize the expanded queries into three types.
 Denote the user input as , and the result of the intent classification as .
-We have,
-When equals to 1, it indicates that, despite supplementing with historical inputs in Section 4.2.1, the user query still lacks critical information.
-Therefore, to ensure the successful execution of the task, we must either automatically retrieve essential information from the knowledge base or request additional details from the user.
-Knowledge Retrieval.
-We utilize the knowledge retrieval function described in Section 4.1.3, using a hybrid approach of Embedding and ElasticSearch to retrieve demonstrations and domain knowledge.
-The demonstrations primarily serve as query clarification prompts.
-We have manually constructed a prompt library of common query clarification forms and stored it along with the necessary knowledge for SiriusBI in the knowledge management module.
+Formally, we classify the query into three categories, as indicated by Eqn.
+(1).
+When , the query is identified as out of the BI scope.
+In this case, the system will politely decline the request and guide the user to ask data analysis-related questions.
+The dialogue analysis process will then restart.
+When , it indicates that, even after supplementing with historical inputs, the user query still lacks key information.
+The system will proactively ask the user to provide the missing information and then restart the analysis process.
+When , the query is complete.
+The system will proceed to the next step of the pipeline.
 Intent Clarification.
-Based on the results of intent recognition and knowledge retrieval, SiriusBI integrates and comprehensively analyzes these key pieces of information during the clarification phase.
-Specifically,
-When is 0, the clarification phase will politely decline and guide the user to ask data analysis-related questions.
-When is 1, the input is likely not reflected in the user’s historical information.
-In this case, the clarification phase will rely on the retrieved information from the knowledge base to enhance the indicators and dimensions.
-It will then ask the user if the supplemented parts meet their expectations.
-If not, the user is prompted to provide the relevant indicators and dimensions in the next round of dialogue.
-This iterative process helps clarify the user’s true intent through follow-up questions, i.e., the true dimensions and indicators of the query.
-When is 2, no further clarification is needed, and the process directly moves to the next module in the SiriusBI pipeline.
-By integrating the intent clarification feature into the dialogue analysis, we have developed the MRD-Q component.
-Through professionally designed clarification prompts, we can understand the user’s true intent more accurately.
-This approach lowers the user’s entry barrier.
-Even if the user provides arbitrary answers, SiriusBI can clarify the user’s correct intent through guided follow-up questions.
-Additionally, it helps to avoid confusion between similar dimensions and indicators, thereby preventing error propagation throughout the entire BI system.
+Although the queries entering this stage are semantically complete, additional processing is required to ensure the generated SQL accurately reflects the user’s intent due to real-world complexities.
+To address this, the module first retrieves domain-specific knowledge from the knowledge base based on the user’s query.
+For instance, the triplet associated with the term “YTD” is defined as: (“YTD”, “Year To Date or Year To Day, referring to the period from the beginning of the current year to the present date.”, “What is the YTD revenue of the mini-program?”).
+Such knowledge enhances the LLM’s understanding of domain-specific terminology.
+Furthermore, the module retrieves task-specific materials, including prompt templates and demonstration examples, to guide the LLM in intent clarification.
+Specifically, the LLM either returns a semantically refined query or interactively presents users with disambiguation options to ensure input precision.
+In summary, we develop the MRD-Q component, which supplements the heuristic dialogue analysis by the intent clarification capability of the LLM.
+Even if the user provides arbitrary answers, SiriusBI can clarify the user’s true intent through guided follow-up questions, therefore improving the successful rate of the task execution.
 3.4. SQL Generation
@@
 In this subsection, we will focus on the implementation approach of the SQL Generation module in SiriusBI, including table selection, as well as the SQL generation methods.
-To address the high cost of model migration across different professional domains, we have also proposed a novel two-step paradigm for SQL generation.
 3.4.1. Table Selection
-Before generating SQL queries, it is necessary to select the appropriate tables from the database based on the user’s input, especially retrieving the relevant table information mentioned in the user’s query.
-As shown in Figure 5, the implementation of table selection is divided into two parts: coarse ranking and reranking.
-In the coarse ranking stage, the number of candidate tables is reduced to a manageable number.
-In the reranking stage, these candidates are re-ranked to identify the most relevant ones for SQL generation.
+Before generating SQL queries, it is necessary to select the appropriate tables from numerous options in an industrial database based on the user’s input.
+This process relies on retrieving the relevant table information mentioned in the user’s query.
+The implementation of table selection is divided into two parts: coarse ranking and re-ranking.
+During the coarse-ranking stage, the number of candidate tables is typically reduced by a factor of ten or more to maintain fewer than 100 candidates, thereby ensuring computational efficiency in the subsequent fine-ranking phase.
+In the re-ranking stage, these candidates are re-ranked to identify the most relevant ones for SQL generation.
 The coarse ranking process consists of two parts: online and offline.
-In the offline part, each data table has a pre-prepared embedding.
-We use the StarRocks (SR) (team, [n.d.]) vector database to store the embeddings of tables to ensure data security.
+In the offline part, each data table is subjected to a private embedding model and transformed into an embedding vector.
+We use the StarRocks (SR) (team, 2023b) vector database to store the embeddings of tables to ensure data security and the embeddings are extracted using the general-purpose pre-trained model m3e-large (Chen et al., 2024c), ensuring high-quality feature representations and domain adaptability.
 SR supports millisecond-level latency retrieval and allows for filtering by specified tags, making it a robust solution for efficient table selection.
@@
 Compared to directly embedding the user’s input, this approach eliminates redundant information, reduces interference from irrelevant information, and increases the recall rate of target tables.
-During reranking, we use different dimensional scoring methods for each table based on the hundreds of tables recalled in the coarse ranking stage. The overall calculation formula is as follows:
+During re-ranking, we use a scoring function balancing multiple metrics for each table based on the hundreds of tables recalled in the coarse ranking stage.
+The overall calculation formula is as follows:
 where represents the semantic recall score during the coarse ranking, which is the cosine similarity score of the vectorized results.
@@
 where represents the similarity score between keyword and table , which is the maximum value of the text similarity with all columns:
-We utilize TF-IDF (Joachims, 1997) as the vectorization method to compute similarity for several reasons: 1) The keywords are relatively short; 2) The calculation needs to be performed in real-time; 3) Gensim-based TF-IDF allows for rapid computation of similarity in batches; 4) Avoid the situation where the vectorized results fail to give proper attention to some keywords in the query when the input is excessively long.
+While TF-IDF (Joachims, 1997) remains effective for real-time matching on short text queries, we perform coarse filtering by restricting candidate tables to those with dynamically adjusted per query complexity.
+3.4.2. Dynamic Strategy Switching Driven by Data Conditions
+In practical BI applications, the data conditions of different clients vary significantly: some clients have abundant labeled data, while others can only provide a limited number of labeled samples due to frequent business changes or data privacy restrictions.
+Similarly, existing SQL generation methods are often tailored to specific scenarios or domains, making migration between private scenarios costly in terms of both deployment effort and data labeling.
+To address this challenge, we propose a data-condition-driven dynamic strategy switching mechanism.
+This mechanism adaptively selects either a one-stage (high-precision fine-tuning) or two-stage (semantic inter- mediate representation) generation strategy based on two key factors: the volume of labeled data and the semantic similarity between domains.
+By doing so, it effectively balances accuracy and cost-efficiency.
+Formally, we define the number of available labeled Query-SQL pairs as the Labeled Data Volume().
+We also quantify the semantic similarity between the target domain and an existing labeled domain (referred to as the source domain) as Domain Similarity()This similarity is computed via the cosine similarity between the embeddings of schema keywords provided by the business side, as follows:
+where and represent the sets of keywords for the target domain and the source domain, respectively (e.g., in the financial domain: ”revenue”, ”quarterly report”).
+is Sentence-BERT (Reimers and Gurevych, 2019) encoding function.
+Based on the aforementioned features, we use rule-based criteria to determine the SQL generation strategy. Specifically:
+where and are hyperparameters that are typically specified based on the actual situation. In our practical experience, is usually set to 500 and is set to 0.7.
 3.4.3. One-Step SQL Generation
-For scenarios where high accuracy in SQL generation is required and the usage domain is relatively stable, we apply a one-step approach that leverages supervised fine-tuning (SFT) of LLMs to allow direct generation of SQL statements from user inputs.
-Notably, we use Hunyuan-70B-SQL as the generation model, which is specifically designed for SQL generation by Tencent.
-While data quality is critical to the accuracy of the fine-tuned model, we develop an automated data pipeline which includes two critical steps: data preparation and data refinement.
-Data Preparation.
-In the data preparation phase, we begin by selecting a batch of natural language requirements from business data, along with their ground-truth SQL generation results.
-Given that these requirements are highly standardized and contain complete contextual information, relying solely on this data poses challenges in ensuring the model’s performance in scenarios with ambiguous information.
-To address this, we first employ a LLM to process each query, preserving the original intent while minimizing the information provided.
-This approach generates a set of semantically ambiguous inputs.
-In addition, we prepare relevant domain knowledge to serve as a reference for these inputs.
-We leverage knowledge retrieval techniques outlined in Section 4.1 to extract supplementary knowledge related to the original queries, while also enabling the LLM to supplement the rules it relies on during the information reduction process.
-An example of data preparation without ground-truth results is as below,
-Finally, the dataset we use for fine-tuning comprises the reduced queries, the supplementary contextual knowledge, and the ground-truth SQL outputs.
-This comprehensive approach enhances the model’s ability to perform effectively in diverse and ambiguous scenarios.
-Data Refinement.
-After fine-tuning on the prepared data, we further introduce an adaptive data refinement method to improve the accuracy of the generation model.
-We develop two strategies to improve the data quality based on error types and user feedbacks.
-The first strategy involves automatically assessing the model’s generated results and categorizing the failed examples into different types, such as time processing errors, unit conversion errors.
-Based on this information, we will prioritize queries that address these needs in the next data preparation phase.
-The second strategy involves conducting small-scale user testing, where users provide simple evaluations of the SQL queries generated by the model.
-Examples that receive low ratings will be annotated and added to the new training set.
-After evaluating the current model with these two strategies, we will iterate on the data preparation phase and fine-tune the model again.
-This process will continue until the model’s accuracy meets the requirements or until we reach the budget limit set by our clients.
+To improve the performance of supervised fine-tuning and reduce annotation efforts, we have developed an automated and iterative data preparation workflow guided by manually labeled data and identified bad cases.
+The overall pipeline is illustrated in Figure 7, consisting of two main modules: data generation and data augmentation.
+Data Generation.
+To perform data generation, we need to utilize the SQL logs from the relevant scenarios and the corresponding schema information of the databases.
+SQL logs are the actual SQL code executed by users, representing the most authentic application data.
+However, since these are directly coded by users in the compiler, we need to use data reverse engineering to capture the users’ true intentions.
+The data reverse engineering functionality is primarily implemented through LLM-based prompts.
+The input information includes SQL, the corresponding schema information, and proprietary domain knowledge base information.
+On the other hand, we generate some question sets using the schema combined with LLM, and then construct some query-SQL pairs through a general SQL generation model.
+The query-SQL pairs obtained through these two methods will enter the data quality evaluation module.
+This module mainly uses the locally deployed Qwen2.5-72B (Yang et al., 2024) model for filtering, combined with manual evaluation to select the final high-quality training set.
+The manually annotated data will serve as the In-Context Learning (Dong et al., 2022) (ICL) demonstrations used in the aforementioned process.
+The specific data generation pipeline can be seen in Figure 7(a).
+Data augmentation.
+This stage enhances data quality through two key approaches: diversity expansion and error-based negative example injection.
+First, we use Self-Instruct (Wang et al., 2023a) with prompts to generate a broader variety of SQL queries from seed data, significantly increasing the dataset’s size and diversity.
+These queries are then evaluated by a LLM (Zhang et al., 2023b), and only high-quality examples are selected for training.
+Second, we introduce negative examples by randomly rewriting cases from the enhanced dataset, based on error categories provided by the client.
+From experience, negative examples typically account for 5% of the training set, helping improve the model’s robustness by teaching it to recognize and avoid common mistakes.
+Together, these strategies ensure a diverse, challenging dataset that enhances the model’s accuracy and reliability.
+Once the data generation and augmentation phases are complete, the refined dataset is used to train the model.
+If necessary, the pipeline can be iterated to further refine the dataset based on additional error analysis or user feedback, ensuring continuous improvement in data quality.
+This iterative refinement continues until the dataset meets the desired standards or until the project reaches the budget constraints specified by our clients.
+Comparison with the previous work on data augmentation for NL2SQL systems.
+Early efforts in this field (Wu et al., 2021; Weir et al., 2020) were generally simple, often relying on basic templates covering only a single table, heuristic methods such as synonym replacement, or traditional deep learning models like RNNs to construct NL-SQL pairs.
+These approaches greatly limited the quality of the generated data.
+More recent work, such as ScienceBenchmark (Zhang et al., 2023a), is closer to our method, proposing an automatic data augmentation system to expand training data.
+The main differences are: (1) Seed Data Source: We use real SQL logs combined with data reverse engineering and NL-to-SQL translation to obtain Query-SQL pairs, while ScienceBenchmark relies on manually created data.
+(2) Augmentation Goals: Besides increasing data diversity, we focus on improving model robustness by injecting noise and negative examples; ScienceBenchmark mainly aims to enlarge dataset size to formalize a high-quality benchmark.
+(3) Augmentation Approach: We directly prompt LLMs to rewrite Query-SQL pairs, whereas ScienceBenchmark uses a multi-step pipeline of template extraction, population, NL question generation, and candidate selection.
 3.4.4. Two-Step SQL Generation
-To meet the demand for a method with strong generalization capabilities and low deployment costs across multiple SQL generation scenarios, we propose a novel two-step generation paradigm.
-This paradigm breaks down the SQL generation task into two stages: the Semantic Model stage and SQL Model stage.
-In the Semantic Model stage, we first structure the user’s input.
-In practical NL2SQL tasks, user inputs are often difficult to control, and incomplete or ambiguous inputs can directly affect the SQL generation accuracy of ChatBI systems.
-This, in turn, reduces user experience and raises the user threshold.
-Therefore, we enhance the semantic information of the inputs before generating SQL from user inputs.
-For a user query , we first perform a retrieval from a knowledge base that includes explanations of domain-specific terms.
-Notably, our knowledge management module (Section 4.1) allows for plug-and-play management, which boosts efficiency of deployment in industrial applications.
-The retrieved domain knowledge, along with , is then processed by a LLM-based Semantic Model for semantic analysis and understanding, resulting in the generation of structured text.
-In our implementation, the structured text takes various forms, such as question rewriting, JSON, and API formats.
-Additionally, during the Semantic Model stage, we employ automated demonstration selection and Chain-of-Thought (CoT) prompting techniques.
-These strategies greatly enhance the accuracy and generalization capability of the Semantic Model for different types of questions.
-In the SQL Model stage, we use another LLM as the SQL generation model, specifically Hunyuan-70B-SQL.
-The enhanced structured information from the Semantic Model serves as the input for the final SQL generation.
-As the structured input provides ample context for SQL generation, the SQL Model can effortlessly generate SQL dialects tailored to various database systems from the same input.
-By structuring and supplementing user questions before generating SQL, the two-step generation paradigm ensures the accuracy of SQL generation, and meanwhile facilitates easy migration across different industrial scenarios.
-This approach significantly reduces the training costs and improves practical deployment efficiency.
-| Category | Example | Description |
-| Descriptive Requests | Which day had the highest revenue? | This category is a general term for basic analysis problems, including correlation, extreme value, difference analysis, etc. |
-| Predictive Requests | Forecast failure rate for the next week. | This category include future trend analysis or label prediction of unseen data. |
-| Diagnostic Requests | Identify anomalies in the sales revenue. | This category includes anomaly detection and metric attribution for numerical data, as well as drill-down for categorical data. |
-| Prescriptive Requests | How to deal with abnormal values of production volume? | This category helps interpret data insights and proposes actionable solutions based on the analysis. |
-| Complex Requests | Forecast the net income for next week and calculate the expected year-over-year change.
-| This category involves executing multiple sub-tasks and there are dependencies between the execution of these sub-tasks.
-|
-4.4. Data Insight
-Data insights are essential to ChatBI systems, as they enable users to extract valuable information from raw data.
-This functionality allows users not only to understand the current state of affairs but also forecast future trends.
-Data insights go beyond traditional descriptive analytics; they reveal hidden patterns in the data and provide deep understanding, including root cause analysis of issues, predictive modeling for the future, and actionable recommendations based on the analysis.
-Such in-depth analysis is crucial for businesses because it helps decision-makers make informed choices, thus driving business growth.
-SiriusBI leverages the capabilities of LLMs to refine the functionalities of data insights, which greatly improves the user experience.
-The data insight module of SiriusBI is designed to address various problems as detailed in Table 1.
-4.4.1. Main Workflow
-The primary workflow for data insights in SiriusBI includes the following steps: automatic data preparation, complex task planning, invocation of insight tools, and generation of insight reports.
-A more detailed process flow will be introduced as follows.
-Automatic Data Preparation.
-The insight tool executes three sequential steps during automatic data preparation: intent rewriting, Text2DSL, and metric-dimension recommendation.
-Initially, intent rewriting categorizes and refines the given insight question into a more specific description.
-Subsequently, Text2DSL converts the question into a Domain-Specific Language (DSL) corresponding to SQL queries.
-By inputting the schema information of the data table along with the insight question, this part outputs structured DSL language that includes essential columns for data insight.
-Finally, the schema of the data table, the DSL output, and the insight question are input into the metric-dimension recommendation part, which suggests additional relevant data columns that could fulfill the insight requirements based on the type of insight task.
-Complex Task Planning.
-The Data Insight module is capable of planning complex tasks, in which a question may contains multiple interdependent requirements.
-It employs an internal planner that decompose these tasks into manageable sub-tasks and iteratively refines the plan based on the execution results of the previous sub-task.
-For instance, consider a scenario where the module is tasked with calculating the month-over-month growth of net new holdings for July 8, 2024, using a business dataset.
-Since the data for that specific date is not available in the data table, the execution process must first include a predictive analysis to estimate the value for July 8 before proceeding with the growth calculation.
-Insight Tool Execution. The capabilities of the data insight module are supported by specialized insight tools. Depending on the specific tasks, the insight tools in SiriusBI primarily include:
-Code Generation.
-For straightforward queries, such as retrieving a specific piece of data, SiriusBI utilizes a LLM to generate the necessary code, which is then executed in a secure sandbox environment.
-Data Modeling.
-For specialized tasks such as prediction and anomaly detection, SiriusBI develops specialized machine learning models and utilizes automated tools for hyperparameter optimization, providing users with accurate output support.
-Visualization.
-For visualization needs, SiriusBI uses a LLM to construct configuration files of visualization settings, which are then used by underlying visualization tools to automatically generate charts.
-Knowledge Retrieval.
-For questions that go beyond the information in the data tables, SiriusBI employs a knowledge retrieval function to extract relevant information (e.g., actions taken in similar scenarios) from the knowledge management module introduced in Section 4.1, thereby assisting in addressing the current task.
-Insight Summary.
-Insight summary is a comprehensive tool to meet various insight requirements.
-It automatically generates data overviews, trend analyses, anomaly diagnostics, and summary recommendations based on the characteristics of the given data table as well as the outputs from previous insight tools.
-Compared with traditional BI systems and the most advanced ChatBI systems, the data insight module proposed in SiriusBI offers users enhanced data analysis capabilities.
-This module supports a wide range of practical application scenarios, enabling users to achieve more usable and reliable data analysis results.
-In addition, it addresses complex and challenging data analysis tasks, meeting the growing demand for dependable analytical solutions.
+To address the need for cross-domain generalization in real-world scenarios and the frequent occurrence of cold-start problems in business applications, we propose a novel two-stage SQL generation paradigm.
+This paradigm divides the SQL generation process into two stages: Semantic Intermediate Representation (SIR) and SQL Generation.
+In the first stage, the construction of SIR focuses on abstracting user intent and data semantics, transforming complex natural language queries into a standardized intermediate form that is easy to handle.
+The rewriting process primarily relies on in-context learning (Dong et al., 2022) to dynamically reconstruct user inputs.
+Additionally, we introduce a Chain-of-Thought (CoT) (Wei et al., 2022) strategy to explicitly guide the model in step-by-step reasoning about the latent intent and constraints of the query, thereby enhancing the quality of the rewritten results.
+Take the query “Analyze consumption of Tencent Cloud BI in 2023” as an example.
+We first extract domain-specific knowledge which includes explanations of terms (“Cloud BI” in this case) and top- demonstrations exhibiting highest similarities with the embeddings of the query as few-shot (Wang et al., 2020b) references.
+The input information in the SIR phase includes field descriptions retrieved from the knowledge base as well as few-shot demonstrations.
+The output of the SIR phase is shown below:
+In the second stage, we use a code-rich LLM for SQL generation.
+The ”Rewriting Query” and the schema information corresponding to the user query together form the prompt.
+By first constructing and refining user queries before generating SQL, this two-step approach ensures high accuracy while enabling seamless migration across diverse industrial scenarios.
 4. Experiments and Results
-To evaluate our framework, in this section, we conduct an in-depth investigation from two insights: 1) On the general NL2SQL task, SiriusBI outperforms the state-of-the-art baselines on both SRD and MRD datasets; 2) The end-to-end design of the SiriusBI system is rational, where the individual performance of each module demonstrates significant superiority.
+To evaluate our framework, in this section, we conduct an in-depth investigation from two insights: 1) On general NL2SQL tasks, SiriusBI outperforms the state-of-the-art baselines on both SRD and MRD datasets; 2) The end-to-end design of SiriusBI is rational, where the individual performance of each module demonstrates significant superiority.
 4.1. Experiment Setups
-Datasets.
-To demonstrate the overall performance of SiriusBI and the effectiveness of its various modules, we conduct comprehensive experiments using datasets from multiple tasks.
+4.1.1. Datasets
+To demonstrate the overall performance of SiriusBI and the effectiveness of its various modules, we conduct comprehensive experiments using both academic benchmarks and industrial-level datasets on multiple tasks.
+We have released our carefully curated academic benchmark, MRD-BIRD.
+Due to data privacy concerns, we provide partially anonymized versions of two industrial-level datasets.
+They are publicly available at https://github.com/TencentBigData-SiriusAI/SiriusBI.
 Detailed descriptions of the datasets used for each task are provided below:
-NL2SQL.
-Although the Spider (Yu et al., 2018) and Bird (Li et al., 2023) datasets are widely used for NL2SQL tasks, their default SRD settings and the gap with industrial scenarios make it difficult to accurately evaluate NL2SQL tasks in actual BI scenarios.
-Therefore, we follow the previous approach (Lian et al., 2024) and select high-quality and reliable data from real application scenarios to construct a muilt-round dialogue (MRD) dataset.
-The dataset contains 55 sets of queries, where each set consists of 2 to 5 rounds.
-In total, the dataset contains 160 queries spanning 35 columns, accompanied by SQL written by human experts for reference.
-Among these queries, 10 sets do not fully include all metrics and dimensions, while another 10 sets exhibit confusion in metrics or dimensions.
-While most baselines are proposed for single-round dialogue (SRD) scenario, to ensure fair comparison, we also supplement the missing information in the MRD dataset and add 22 more single-round queries, thereby creating a SRD dataset with totally 182 queries for evaluation purposes.
+SRD NL2SQL.
+To validate the effectiveness of the SQL generation module, we utilize the BIRD (Li et al., 2023), a challenging large-scale database NL2SQL evaluation benchmark aimed at narrowing the gap between academic research and practical applications.
+It includes 95 large databases and high-quality (user question, SQL) pairs covering 37 professional fields.
+Besides, we collect 182 SRD queries spanning 10669 tables and involving 43 columns from 5 different business scenarios, thereby creating a SRD dataset termed as SRD-Industry.
+MRD NL2SQL.
+This task better simulates real-world BI scenarios by evaluating both intent understanding and SQL generation in multi-round dialogue (MRD) settings.
+Due to the lack of a standard MRD NL2SQL benchmark, we expand part of the SRD dialogues from the BIRD dataset into 3–5 round dialogues.
+Specifically, we use an LLM to decompose complex SQL into turn-level queries, then generate corresponding user questions that reflect natural MRD conversation style—building on prior context without repetition.
+We also create MRD-Industry, a new dataset matching the database scale of SRD-Industry, with 55 query sets spanning 2–5 rounds each (160 queries over 35 columns).
+To increase complexity, we select 10 sets and remove certain metrics and dimensions, and select another 10 sets in which we inject confusion by replacing metrics or dimensions with ambiguous alternatives.
 Knowledge Management.
 To evaluate the knowledge management module, we focus on two main aspects: 1) Assessing its impact on NL2SQL tasks; 2) Evaluating the performance of knowledge retrieval.
-In the latter evaluation, we utilize the knowledge base corresponding to the MRD dataset, which contains knowledge across six dimensions: table, column, value, term, udf, and alias, totaling 7704 pieces of knowledge.
+In the latter evaluation, we utilize the knowledge base corresponding to the MRD-Industry dataset, which contains knowledge across six dimensions: table, column, value, term, udf, and alias, totaling 7704 pieces of knowledge.
 Data Insight.
@@
 Note that, other types of requests are not included in the dataset, as it’s difficult to evaluate the correctness of answers to open-ended questions.
-Evaluation Metrics.
-In NL2SQL tasks, we follow previous work (Lian et al., 2024) and applies the Useful Execution Accuracy (UEX) metric.
-While other metrics like Execution Accuracy and Exact Match require the generated query matches the ground-truth query precisely, UEX recognizes a query as correct if the execution results of the generated query align with the actual results.
-For tasks beyond NL2SQL within SiriusBI, such as knowledge retrieval in knowledge management, comprehensive analysis in data insight, we employ the retrieval recall rate and the execution accuracy rate, respectively.
-Baselines. We compare the NL2SQL capability of SiriusBI with three state-of-the-art baselines as below:
-DIN-SQL (Pourreza and Rafiei, 2023) breaks down the NL2SQL task into discrete sub-tasks, providing specific prompts for each sub-task to GPT-4o through the completion of each component.
-MAC-SQL (Wang et al., 2023) employs a multi-agent framework to handle NL2SQL tasks, involving three main agents: selector, decomposer, and refiner.
-These agents work together to enhance the accuracy and efficiency of SQL generation.
-MRD-SQL is a reproduction of previous work (Lian et al., 2024).
-We reproduce the basic MRD system integrated with an SQL generation model.
-To ensure fair comparison, we use the same base LLM as in SiriusBI.
-SiriusBI employs all modules introduced in this paper.
-For evaluation of NL2SQL, the data insight module is excluded.
-SiriusBI processes user inputs through dialogue analysis and intention querying before generating SQL queries.
-We use “one-step” and “two-step” to distinguish between the different SQL generation methods in SiriusBI.
-LLM Settings.
-In SiriusBI and MRD-SQL, we utilize the Hunyuan-70B-SQL, specifically designed for SQL scenarios, developed by Tencent’s Hunyuan LLM group.
-For other baselines, we employ the most advanced OpenAI API (Ope, [n.d.]) as the foundational LLM, specifically the ”gpt-4o-2024-08-06” version, to ensure optimal performance in all compared baselines.
-Training Setting.
-For LLM fine-tuning in one-step SQL generation method, we use four nodes equipped with a total of 32 NVIDIA-A100-80GB GPUs, setting the training to 3 epochs, a learning rate of 3e-5, a batch size of 2, and a sequence length of 8192 tokens.
-The fine-tuning is conducted on a private dataset consisting of over 9,000 entries collected from industrial scenarios.
-5.2. Experiment with NL2SQL
-In this experiment, we primarily evaluate the performance of various baselines on the SRD and MRD datasets.
-Note that, as an analysis module following NL2SQL, the data insights module do not participate in this part.
-5.2.1. Results on SRD Dataset
-We first conduct experiments on SRD dataset and present the results in Table 2.
-Our findings indicate that SiriusBI with both two SQL generation methods show a clear improvement over other baselines on the dataset.
-Specifically, when compared to competitive baselines such as GPT-4o and its derivatives (DIN-SQL and MAC-SQL), SiriusBI achieves an improvement of at least 40% in UEX.
-In comparison to the SOTA ChatBI implementation, MRD-SQL, the two-step and one-step variants of SiriusBI achieve improvements of 17.34% and 21.39%, respectively, which highlights the effectiveness of SiriusBI NL2SQL pipeline.
-Interestingly, there is a relatively slight performance difference between two variants of SiriusBI.
-Concretely, the one-step method, which has been carefully fine-tuned on domain-specific data, gets 4.04% higher in UEX than the SFT-free two-step method.
-However, it is worth noting that the data for SFT took two weeks of annotation by four experts, and SFT itself also took 12 hours.
-Therefore, we recommend that users seeking higher accuracy use the one-step approach, while those prioritizing faster deployment efficiency should opt for the two-step approach.
-| Baseline | UEX |
-| DIN-SQL | 30.06% |
-| MAC-SQL | 36.42% |
-| MRD-SQL | 61.85% |
-| SiriusBI two-step | 79.19% |
-| SiriusBI one-step | 83.24% |
-5.2.2. Results on MRD Dataset
-Next, we evaluate SiriusBI on the MRD dataset and present the specific experimental results in Table 4.
-Our findings reveal that single-round methods perform poorly on the MRD dataset, where the UEX falls below 16%.
-In contrast, the two-step and one-step variants of SiriusBI achieve UEX scores of 55.00% and 62.50%, respectively, demonstrating the superiority of SiriusBI in real-world NL2SQL tasks when compared to the state-of-the-art multi-round design, MRD-SQL.
-Remarkably, we also investigate the impact of SFT on the MRD dataset.
-By comparing the performance of SiriusBI one-step and MRD-SQL with and without fine-tuning, we observe that SFT yields significant performance gains at the cost of higher resource investment.
-Furthermore, we find that though SiriusBI two-step does not incorporate SFT, it still improves SiriusBI one-step without SFT by 5.62% through effective retrieval of domain knowledge.
-This leads to a similar conclusion that users can choose the one-step method for higher accuracy and the two-step method for faster deployment.
-5.2.3. In-production Results.
-Besides experiments on datasets collected from industrial scenarios, we also evaluate SiriusBI in production.
-So far, SiriusBI has been deployed in Tencent’s finance, advertising and cloud sectors.
-For Tencent Finance and Advertisement, we apply the one-step method by fine-tuning LLMs on elaborately collected domain-specific datasets, achieving in-production accuracies are 97% and 89%, respectively.
-In the case of Tencent Cloud where SiriusBI is available to business clients, we apply the two-step method to reduce deployment costs for users.
-The average accuracy across 14 clients are 91%.
-Interestingly, we find that the in-production results are higher than those reported on the SRD and MRD datasets.
-We attribute this to the differences in problem distribution.
-In industrial scenarios, users’ questions may be simpler, while the datasets include high-quality and complex questions, which are more suitable for assessing the system’s specialized capabilities.
-| Baseline | UEX |
-| DIN-SQL | 11.25% |
-| MAC-SQL | 15.63% |
-| MRD-SQL w/o SFT | 22.50% |
-| MRD-SQL | 28.13% |
-| SiriusBI two-step | 55.00% |
-| SiriusBI one-step w/o SFT | 49.38% |
-| SiriusBI one-step | 62.50% |
-5.3. Systematic Deep Analysis Experiments
-In addition to NL2SQL, we also conduct a deeper analysis of the effectiveness of various components in SiriusBI.
-In this section, we primarily focus on analyzing modules related with MRD-Q, knowledge management and data insights.
+4.1.2. Evaluation Metrics
+For NL2SQL tasks on the BIRD dataset, we adhere to the standard evaluation metrics as outlined in (Li et al., 2023), utilizing Execution Accuracy (EX) and Valid Efficiency Score (VES).
+For NL2SQL tasks on our internal datasets, we follow previous works (Lian et al., 2024; Floratou et al., 2024) and employ the Useful Execution Accuracy (UEX) metric, which considers a query correct if its intent aligns with that of the golden query.
+This makes UEX more suitable for business intelligence scenarios.
+4.1.3. Baselines and Model Configurations
+We compare the NL2SQL capability of SiriusBI with three state-of-the-art baselines as below: (1) DIN-SQL (Pourreza and Rafiei, 2023); (2) MAC-SQL (Wang et al., 2023b); (3) MRD-SQL (Lian et al., 2024); (4) SiriusBI.
+For the baseline models, we followed their publicly available usage instructions as closely as possible, including base models and prompt designs.
+For DIN-SQL and MAC-SQL, we used their original configurations and prompts without modification.
+We reproduce the NL2BI module of the previous work (Lian et al., 2024) as MRD-SQL.
+In our method and MRD-SQL, we use Qwen2.5-Coder-32B (Hui et al., 2024) as the base model for two-step SQL generation on SRD and MRD datasets.
+For one-step SQL generation on the BIRD dataset (and MRD-BIRD), we use XiYanSQL-QwenCoder-32B (Gao et al., 2024).
+This choice is motivated by its demonstrated state-of-the-art performance on this benchmark, which obviates the need for additional fine-tuning.
+For industrial datasets, we fine-tuned Qwen2.5-Coder-32B to better fit domain data.
+Specifically, we use four nodes equipped with a total of 32 NVIDIA A100-40GB GPUs, setting the training to 3 epochs, a learning rate of 3e-5, a batch size of 2, and a sequence length of 8192 tokens.
+The fine-tuning dataset comprises over 9,000 entries collected from real-world industrial scenarios.
+Overall, we ensured fair comparison by aligning baseline setups with their original methods and optimizing prompts within their intended use.
+4.2. End-to-End NL2SQL Performance
+In this experiment, we evaluate the end-to-end NL2SQL performance of various baselines on the SRD and MRD datasets.
+Note that, as an analysis module following NL2SQL, the data insight module is not included in this part.
+4.2.1. Results on SRD Datasets
+We conduct experiments on two SRD datasets—BIRD’s dev set and SRD-Industry—with results summarized in Table 1.
+Our method, SiriusBI, employing both one-step and two-step SQL generation, consistently outperforms baseline approaches.
+The improvement is especially pronounced on the SRD-Industry dataset, where SiriusBI achieves at least a 40% increase in UEX.
+This significant gain is attributed to SiriusBI ’s explicit design to tackle real-world challenges such as ambiguous fields and wide tables, which general-purpose models often struggle with.
+We exclude MRD-SQL from this comparison since it differs from SiriusBI only in the multi-round dialogue analysis module, leading to identical performance on SRD datasets.
+Additionally, the one-step variant of SiriusBI, fine-tuned on domain-specific data, surpasses the two-step method by 4.05% in UEX.
+This suggests that supervised fine-tuning provides a measurable advantage in handling domain-specific queries, further enhancing its performance.
+Methods | BIRD | SRD-Industry | |
+EX(%) | VES(%) | UEX(%) | |
+DIN-SQL + GPT-4o | 50.72 | 58.79 | 30.06 |
+MAC-SQL + GPT-4o | 57.56 | 58.04 | 36.42 |
+SiriusBI two-step | 65.32 | 67.87 | 79.19 |
+SiriusBI one-step | 68.97 | 70.89 | 83.24 |
+4.2.2. Results on MRD Datasets
+Next, we evaluate SiriusBI on two MRD datasets: MRD-BIRD and MRD-Industry, and present the experimental results in Table 2.
+Our findings reveal that NL2SQL methods designed for single-round dialogue perform poorly on MRD datasets.
+Specifically, the UEX scores of DIN-SQL and MAC-SQL on the MRD-Industry dataset fall below 16%, highlighting their limitations in handling multi-round dialogue scenarios.
+In contrast, the two-step and one-step variants of SiriusBI achieve UEX scores of 55.00% and 62.50%, respectively, demonstrating the consistent superiority compared to MRD-SQL in real-world NL2SQL tasks.
+Methods | MRD-BIRD EX(%) | MRD-Industry UEX(%) |
+DIN-SQL + GPT-4o | 41.69 | 11.25 |
+MAC-SQL + GPT-4o | 42.35 | 15.63 |
+MRD-SQL two-step | 46.98 | 22.50 |
+MRD-SQL one-step | 50.61 | 28.13 |
+SiriusBI two-step | 48.03 | 55.00 |
+SiriusBI one-step | 51.14 | 62.50 |
+It is worth noting that on the MRD-BIRD dataset, the performance gap between SiriusBI and MRD-SQL is narrower.
+Specifically, the one-step variant of MRD-SQL achieves a higher EX score than the two-step variant of SiriusBI.
+We attribute this phenomenon to the absence of user interaction for intent clarification in the experiments on MRD-BIRD, which diminishes the advantage of SiriusBI.
+However, by incorporating domain-specific knowledge during intent clarification, SiriusBI still outperforms MRD-SQL when utilizing the same SQL generation method, underscoring the effectiveness of our approach in leveraging domain-specific insights to enhance performance.
+4.2.3. In-production Results
+In addition to experiments on industrial datasets, we evaluate SiriusBI in production across multiple Tencent business domains.
+Before deployment, we apply the proposed data-condition-driven dynamic strategy switching mechanism to select the most suitable SQL generation approach based on each domain’s data characteristics.
+For Tencent Finance and Advertising, clients provide sufficient labeled data () and the source and target domains are identical, resulting in a semantic similarity .
+Under these conditions, the mechanism selects the one-step fine-tuning strategy, achieving in-production accuracies of 97% and 93%, respectively.
+Conversely, Tencent Cloud serves diverse users with limited labeled data and low domain similarity between our available labeled data.
+Thus, the mechanism chooses the two-step strategy, which attains an average accuracy of 96% across 14 randomly selected clients.
+4.2.4. Discussion
+We observe that the in-production results are higher than those reported on the SRD and MRD datasets.
+This is mainly due to three factors: 1) Knowledge Completeness: Although (MRD-)BIRD datasets provide extra knowledge, their coverage is limited compared to our comprehensive internal knowledge base, which better supports accurate SQL generation in real scenarios.
+2) Data Distribution and Query Complexity: Real-world user queries in specific applications tend to be more uniform and simpler, allowing fine-tuned models to achieve higher accuracy.
+In contrast, SRD and MRD contain diverse and complex questions from multiple domains, making them more challenging benchmarks.
+3) Iterative Feedback and Model Refinement: In-production models benefit from multiple rounds of feedback and continuous improvement guided by real failure cases (see Section 3.4.3).
+SRD and MRD results, however, reflect single-cycle training without such iterative updates, limiting their performance.
+In summary, richer knowledge, simpler real-world queries, and ongoing model refinement explain the higher accuracy observed in production compared to benchmark datasets.
+4.3. Detailed Analysis
 4.3.1. Evaluation of MRD-Q
-In this part, we present the ablation study on dialogue analysis and intention querying using the MRD dataset/ The experimental results are displayed in Table 4.
+In this part, we present the ablation study on dialogue analysis and intention querying using the MRD-Industry dataset.
+The experimental results are displayed in Table 3.
 Based on the experimental results, we can observe that, both the SiriusBI one-step and SiriusBI two-step heavily rely on the dialogue analysis and the intention querying modules.
-Specifically, when both modules are removed (without MRD-Q), the performance of SiriusBI declines significantly, with drops of 34.37% and 35.62% for one-step and two-step designs, respectively.
-Similarly, the absence of the intention querying module results in a performance degradation of 11.25% and 21.25% for SiriusBI.
+Specifically, when both modules are removed (without MRD-Q), the performance of SiriusBI declines significantly, with drops of 15.62% and 34.38% for one-step and two-step designs, respectively.
+Similarly, the absence of the intention querying module results in a performance degradation of 6.25% and 21.25% for SiriusBI.
 This indicates that the dialogue analysis and intention querying modules are critical to the overall effectiveness of SiriusBI.
-In addition, the ablation study reveals that SiriusBI one-step fine-tuned on a private dataset exhibits greater robustness compared to SiriusBI two-step.
-Concretely, the impact of removing intention querying on SiriusBI one-step is less pronounced, resulting in a performance drop of only 11.25%, compared to a more significant decline of 21.25% for SiriusBI two-step.
-The difference can be attributed to the fact that SiriusBI one-step acquires domain-specific knowledge from the training data during the learning process, which helps to mitigate the negative effects of lacking conversational understanding and intention querying.
 Baseline | UEX | |
@@
 SiriusBI one-step | 62.50% | - |
+In addition, the ablation study reveals that SiriusBI one-step fine-tuned on a private dataset exhibits greater robustness compared to SiriusBI two-step.
+Concretely, the impact of removing intention querying in SiriusBI one-step is less pronounced, resulting in a performance drop of only 6.25%, compared to a more significant decline of 21.25% for SiriusBI two-step.
+The difference can be attributed to the fact that SiriusBI one-step acquires domain-specific knowledge from the training data during the learning process, which helps to mitigate the negative effects of lacking conversational understanding and intention querying.
+4.3.2. Evaluation of Table Selection
+While table selection directly impacts the UEX of SQL selection, we conduct an ablation study on this component.
+We use the queries in the SRD dataset and evaluate the performance using the Recall@5 metric as described before.
+The experimental results are shown in Table 4.
+Baseline | Recall@5 | |
+SiriusBI w/o Embed & Heat | 77.47% | -13.74% |
+SiriusBI w/o Heat | 79.12% | -12.01% |
+SiriusBI w/o Embed | 83.52% | -7.69% |
+SiriusBI | 91.21% | - |
+It is clear that both the embedding and heat information enhance the recall performance of table selection, which indicates the effectiveness of the combined ranking function in Equation 2.
 4.3.3. Evaluation of Knowledge Management
@@
 These results highlight the critical role of the knowledge management module in enhancing the overall performance of NL2SQL, as both intention querying and SQL generation heavily depend on the availability of relevant knowledge.
+Next, we evaluate the module using a dataset for knowledge retrieval with 7704 samples.
+In this evaluation, we compare SiriusBI against two widely used baseline models: BM25 (Robertson and Sparck Jones, 1976) and Vector (Karpukhin et al., 2020), both of which are prominent in the field of text retrieval.
+We assess the performance of these methods using two key metrics: KR Recall measuring the recall rate for retrieving relevant knowledge entries across all keywords and SL Recall measuring the recall rate specifically for linking table columns.
+The experimental results demonstrate the following performance: BM25 achieves a KR Recall of 90.28% and an SL Recall of 89.58%; Vector achieves a KR Recall of 93.06% and an SL Recall of 93.06%; SiriusBI achieves a KR Recall of 95.14% and an SL Recall of 93.75%.
+SiriusBI outperforms the baselines across both metrics, thereby ensuring that other modules leveraging knowledge management within SiriusBI can access accurate and relevant additional information.
 Baseline | UEX | |
@@
 SiriusBI one-step | 62.50% | - |
-Next, we evaluate the module using the dataset for knowledge retrieval with 7704 samples.
-In this evaluation, we compare SiriusBI against widely used baseline models BM25 (Robertson and Sparck Jones, 1976) and Vector (Karpukhin et al., 2020), both of which are prominent in the field of text retrieval.
-We assess the performance of these methods using the recall rates of knowledge retrieval for all keywords (KR) and schema linking (SL) for table columns only.
-Our results indicate that SiriusBI outperforms the baselines across both metrics, thereby ensuring that other modules leveraging knowledge management within SiriusBI can access accurate and relevant additional information.
-| Baseline | KR Recall | SL Recall |
-| BM25 | 90.28% | 89.58% |
-| Vector | 93.06% | 93.06% |
-| SiriusBI | 95.14% | 93.75% |
 4.3.4. Data Insight Evaluations
-Finally, we conduct experiments on the Data Insight module. Remind that the data insight dataset contains quantifiable questions involving comprehensive analysis and complex task planning.
-| Baseline | Accuracy |
-| Command-R+ | 75% |
-| Qwen2.5 | 80% |
-| GPT-4o | 83% |
-| SiriusBI | 95% |
-For Comprehensive Analysis tasks, we use 100 real-world queries from the financial domain and conducts comparative experiments with three powerful LLM baselines: GPT-4o (Ope, [n.d.]), Qwen2.5 (Team, 2024) and Commnd R+ (cohere.ai Team, [n.d.]).
-The accuracy of responses to these queries serves as the evaluation criterion, and to ensure the reliability of the experimental results, each query is tested for three times.
-The specific experimental results are shown in Table 7.
-The findings indicate that the Data Insight module of SiriusBI maintains a notable performance advantage even against the currently most effective LLMs.
-The function call architecture (Kim et al., 2024) in SiriusBI enables it to outperform GPT-4o by 12% in Comprehensive Analysis tasks.
-For Complex Task Planning, we compare the Data Insight module with a method that only performs tool invocation on 50 high-quality queries.
-Compared to the simple tool invocation method, the multi-agent architecture adopted by SiriusBI shows a significant advantage, by improving the accuracy from 70% to 100% on this dataset.
-To facilitate understanding of the execution process of Complex Task Planning, we provide an example as follows:
-6. Related Work
-The integration of Artificial Intelligence (AI) and Machine Learning (ML) has fundamentally transformed data-driven decision-making and strategic planning within the realm of Business Intelligence (BI) (Rane et al., 2024).
-Machine Learning and Predictive Analytics are essential for enhancing BI, providing advanced tools for forecasting and identifying patterns in extensive datasets (Lee et al., 2022; Tamang et al., 2021).
-Additionally, Natural Language Processing (NLP) and Text Analytics enable the extraction of valuable insights from unstructured data, empowering businesses to understand customer sentiment, automate customer service, and enhance decision-making processes (Kaur et al., 2019; Ittoo et al., 2016).
-By leveraging the Internet of Things (IoT) and Big Data, businesses can increase their resilience and adaptability in an ever-evolving market (Alghamdi and Al-Baity, 2022; Klisarova-Belcheva et al., 2017).
-Advanced AI models, such as ChatGPT (Ope, [n.d.]), play a significantly augment BI by providing intelligent data interpretations and improving human-machine interactions, making data more accessible and comprehensible for non-technical users (Linkon et al., 2024).
-A BI system typically stores vast amounts of information, such as customer and product sales data.
-In traditional BI systems, users need to use SQL to aggregate the required data and perform downstream analysis.
-Therefore, to lower the barrier to using BI capabilities and to create an intelligent BI system, NL2SQL has become a core task.
-Recently, Large Language Models (LLMs), including ChatGPT (Ope, [n.d.]), Codex (cod, [n.d.]) and Claude 3 (cla, [n.d.]), have demonstrated remarkable potential in addressing NL2SQL tasks.
-These advancements benefit from the introduction of prompt engineering (Giray, 2023), in-context learning (Dong et al., 2022), and the chain-of-thought approach (Wei et al., 2022).
-Research into prompt-based NL2SQL methods has not only provided new technical paradigms for addressing the challenges of this task but has also revealed opportunities for reducing technical costs in industrial applications.
-For instance, methods such as DIN-SQL (Pourreza and Rafiei, 2023), C3 (Dong et al., 2023), MAC-SQL (Wang et al., 2023), and SQL-PaLM (Sun et al., 2023), designed for models like ChatGPT and PaLM (Narang and Chowdhery, 2022), have demonstrated significant improvements in the performance on NL2SQL tasks.
-Prior to the emergence of LLMs, solutions to the NL2SQL problem typically involved training an ”encoder-decoder” model in a sequence-to-sequence format (Cai et al., 2021; Cao et al., 2021; Choi et al., 2021; Gan et al., 2021; Wang et al., 2020), which heavily relied on a substantial amount of database and SQL-related data.
-The rise of large models has introduced new technical paradigms for traditional train-based methods.
-For example, Codes (Li et al., 2024) propose a training method tailored for the NL2SQL task, which enhances the performance of LLMs in this area.
-However, to achieve better performance in the NL2SQL task, some studies (Maamari et al., 2024; Pourreza et al., 2024; Caferoğlu and Ulusoy, 2024) opt to train LLMs such as ChatGPT (Ope, [n.d.]) and Gemini (gem, [n.d.]), which often comes with a high cost in practical applications.
-In summary, although there have been pioneering efforts in academia to integrate LLMs into BI systems, these works have not fully addressed practical needs.
-From a practical perspective, we must also consider the flexible user interaction requirements, efficient NL2SQL capabilities, and a complete BI workflow.
-Therefore, how to incorporate LLMs into BI and develop a new generation of intelligent BI tools remains an open question.
+Finally, we conduct experiments on the Data Insight module. Remind that the data insight dataset contains 80 quantifiable questions involving comprehensive analysis and complex task planning.
+For evaluation, we use real-world queries from the financial domain and conduct comparative experiments with three powerful LLM baselines: GPT-4o (team, 2022), Qwen2.5 (Hui et al., 2024) and Commnd R+ (cohere.ai Team, 2024).
+The accuracy of responses to these queries serves as the evaluation criterion, and to ensure the reliability of the experimental results, each query is tested three times.
+The results demonstrate that the Data Insight module of SiriusBI achieves an accuracy of 95%, significantly outperforming the baseline models.
+Specifically, Command-R+ achieves an accuracy of 75%, Qwen2.5 achieves 80%, and GPT-4o achieves 83%.
+These findings indicate that the Data Insight module of SiriusBI maintains a notable performance advantage even against the most effective LLMs currently in use.
+This superior performance can be attributed to the function call architecture in SiriusBI, which enables it to outperform GPT-4o by 12% in comprehensive data analysis tasks.
+Specifically, for complex task planning, we compare the Data Insight module with a method that only performs tool invocation on the aforementioned high-quality queries.
+Compared to the simple tool invocation method, the MRD-agent architecture adopted by SiriusBI shows a significant advantage, improving the accuracy from 70% to 100% on this dataset.
+4.4. Case Study
+In this part, we use an exemplar to demonstrate the whole workflow through a BI interaction.
+At the beginning, the user raises a question – What is the income of the Company A in 2024?
+The Dialogue Analysis module evaluates the semantic integrity of this query.
+In fact, this query contains both a “dimension” (2024) and a “metric” (income), making it semantically complete.
+The query then proceeds to the second stage.
+The Intention Querying module recalls relevant information from the knowledge base based on the input query.
+Specifically, a piece of retrieved knowledge is a triplet – (company, by default, without special instructions, refers to Tencent, “What is gross profit of the company in 2023?”), which provides an explanation for the term “company”.
+The Intention Querying module further prompts the user for clarification as there still exists ambiguity within the query.
+In this case, the word “income” is ambiguous (with tax or not), the user will be asked whether it refers to the column “shouldincome” or “shouldincome_after”.
+Afterwards, the SQL generation module selects the appropriate tables and translates the user query into an SQL query.
+The generated SQL statement is SELECT SUM(shouldincome_after) AS total_income FROM revenue_by_quarter WHERE YEAR
+(ftime) = 2024 AND cname = Company A.
+The Data Insight module receives the SQL queries and their corresponding execution results.
+Based on the user’s additional insight requirements like “Perform dimension attribution analysis on the revenue of Company A in 2024”, the module analyzes the user’s intent, decomposes tasks through the Planner, performs data preparation and invokes attribution tools, and finally generates an insight report, as shown in Figure 8.
+4.5. Utility and Usability Study
+To evaluate whether SiriusBI effectively improves users’ productivity and experience in real-world scenarios, we conducted a gray-release test within a business analysis team at Tencent data platform.
+This study aimed to quantify the system’s impact on operational efficiency and gather subjective feedback on usability.
+We randomly divided 40 business analysts into two groups – Group A (20 analysts): granted access to SiriusBI for daily tasks; Group B (20 analysts): continued using legacy tools.
+The experiment lasted for two weeks, during which we monitored both groups’ workflows.
+To ensure comparability, we controlled that tasks were distributed evenly across groups based on complexity and historical completion times.
+We find that Group A exhibited a 30.2% reduction in average task completion time compared to Group B.
+This efficiency gain is attributed to SiriusBI’s integrated workflow and advanced technical design.
+What’s more, Group A provided an average SUS score of 84.2/100 (SD = 6.5), categorizing SiriusBI as “excellent” according to the SUS benchmark (Brooke et al., 1996).
+Key strengths highlighted in feedback included: (1) 89% of users reported that domain-specific knowledge grounding reduced manual schema lookup time.
+(2) 92% agreed that SiriusBI’s data insight module “provided reliable support for business decision-making” (e.g., identifying root causes of revenue fluctuations).
+The gray-release test confirms that SiriusBI enhances both productivity and user experience by addressing the functionality deficiencies.
 5. Conclusion
-In this paper, we propose SiriusBI, an deployed end-to-end LLM-based Business Intelligence system specifically designed to address three key challenges encountered in industrial scenarios.
-Our approach includes: 1) implementing multi-round dialogue with querying capabilities to overcome interaction limitations; 2) employing two SQL generation strategies to alleviate performance bottlenecks; and 3) establishing a comprehensive end-to-end workflow of the entire BI process to address functionality deficiencies.
-Currently, SiriusBI is deployed across Tencent’s finance, advertising and cloud sectors, serving dozens of enterprise clients.
-Extensive experiments on industrial datasets demonstrate that SiriusBI clearly outperforms competitive LLM-based baselines in both SQL generation and data insights generation.
+In this paper, we propose SiriusBI, a LLM-powered solution for business intelligence, specifically designed to address three key challenges encountered in industrial business intelligence scenarios.
+Our system 1) establishes a comprehensive end-to-end workflow of the entire business intelligence process to address functionality deficiencies; 2) implements multi-round dialogue with querying capabilities to overcome interaction limitations; 3) adopts a data-conditioned SQL generation method selection strategy to alleviate the great cost brought by cross-domain migration in SQL generation.
+Currently, SiriusBI is deployed across Tencent’s finance, advertising, and cloud sectors, serving enterprise clients with various profiles.
+Extensive experiments on public benchmark and industrial datasets demonstrate that SiriusBI clearly outperforms competitive LLM-based baselines in both SRD and MRD-based NL2SQL tasks.
+Detailed analysis on different modules demonstrate the rationality of our design.
+User studies further confirm that SiriusBI enhances both productivity and user experience.
 Acknowledgements.
-We thank the researchers and engineers from the Data Platform department, Technology and Engineering Group (TEG) of Tencent for their technical and engineering support.References
+We thank the engineers from the Data Platform department, Technology and Engineering Group (TEG) of Tencent for their technical and engineering support.References
 - (1)
-- Apa ([n.d.]) [n.d.]. Apache Superset. https://superset.apache.org/
-- cla ([n.d.]) [n.d.]. Claude3. https://anthropic.com/claude
-- gem ([n.d.]) [n.d.]. Gemini. https://gemini.google.com/
-- Ope ([n.d.]) [n.d.]. Introduction of OpenAI Text Generation APIs. https://platform.openai.com/docs/guides/text-generation
-- cod ([n.d.]) [n.d.]. OpenAI CodeX. https://openai.com/index/openai-codex
-- Alghamdi and Al-Baity (2022) Noorah A. Alghamdi and Heyam H. Al-Baity. 2022. Augmented Analytics Driven by AI: A Digital Transformation beyond Business Intelligence. Sensors 22, 20 (2022), 8071.
+- Affolter et al. (2019) Katrin Affolter, Kurt Stockinger, and Abraham Bernstein. 2019. A comparative survey of recent natural language interfaces for databases. The VLDB Journal (2019).
+- Bhagwan et al.
+(2014) Ranjita Bhagwan, Rahul Kumar, Ramachandran Ramjee, George Varghese, Surjyakanta Mohapatra, Hemanth Manoharan, and Piyush Shah.
+2014.
+Adtributor: Revenue debugging in advertising systems.
+In 11th USENIX Symposium on Networked Systems Design and Implementation (NSDI 14).
+- Box et al. (2015) George EP Box, Gwilym M Jenkins, Gregory C Reinsel, and Greta M Ljung. 2015. Time series analysis: forecasting and control. John Wiley & Sons.
+- Brooke et al. (1996) John Brooke et al. 1996. SUS-A quick and dirty usability scale. Usability evaluation in industry (1996).
 - Caferoğlu and Ulusoy (2024) Hasan Alp Caferoğlu and Özgür Ulusoy. 2024. E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL. arXiv preprint arXiv:2409.16751 (2024).
@@
 2541–2555.
+- Chaudhuri et al. (2011) Surajit Chaudhuri, Umeshwar Dayal, and Vivek Narasayya. 2011. An overview of business intelligence technology. Commun. ACM (2011).
 - Chen et al.
@@
 - Chen et al.
+(2024c) Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian, and Zheng Liu.
+2024c.
+Bge m3-embedding: Multi-lingual, multi-functionality, multi-granularity text embeddings through self-knowledge distillation.
+arXiv preprint arXiv:2402.03216 (2024).
+- Chen et al.
 (2024a) Xiaoxin Chen, Meng Wu, and Mangning Wang.
@@
 Application of business intelligence under deep neural network in credit scoring of bank users.
-Journal of Computational Methods in Sciences and Engineering 24, 3 (2024), 1585–1604.
+Journal of Computational Methods in Sciences and Engineering (2024).
+- Cheng et al.
+(2009) Hilary Cheng, Yi-Chuan Lu, and Calvin Sheu.
+2009.
+An ontology-based business intelligence application in a financial knowledge management system.
+Expert Systems with Applications (2009).
 - Choi et al.
@@
 Computational Linguistics 47, 2 (2021), 309–332.
-- Choudhury et al.
-(2023) Muntabir Hasan Choudhury, Lamia Salsabil, Himarsha R Jayanetti, Jian Wu, William A Ingram, and Edward A Fox.
-2023.
-MetaEnhance: Metadata Quality Improvement for Electronic Theses and Dissertations of University Libraries.
-In 2023 ACM/IEEE Joint Conference on Digital Libraries.
-61–65.
-- cohere.ai Team ([n.d.]) cohere.ai Team. [n.d.]. The Cohere Platform. https://docs.cohere.com
+- cohere.ai Team (2024) cohere.ai Team. 2024. The Cohere Platform. https://docs.cohere.com
+- Devlin et al.
+(2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
+2019.
+Bert: Pre-training of deep bidirectional transformers for language understanding.
+In Proceedings of the 2019 conference of the North American chapter of the association for computational linguistics: human language technologies, volume 1 (long and short papers).
 - Dong et al.
@@
 - Elasticsearch (2018) BV Elasticsearch. 2018. Elasticsearch. software], version 6, 1 (2018).
+- Floratou et al.
+(2024) Avrilia Floratou, Fotis Psallidas, Fuheng Zhao, Shaleen Deep, Gunther Hagleither, Wangda Tan, Joyce Cahoon, Rana Alotaibi, Jordan Henkel, Abhik Singla, et al.
+2024.
+NL2SQL is a solved problem… Not!.
+In CIDR.
 - Gan et al.
@@
 Patterns 4, 10 (2023), 100839.
+- Gao et al.
+(2024) Yingqi Gao, Yifu Liu, Xiaoxia Li, Xiaorong Shi, Yin Zhu, Yiming Wang, Shiqi Li, Wei Li, Yuntao Hong, Zhiling Luo, Jinyang Gao, Liyu Mou, and Yu Li.
+2024.
+XiYan-SQL: A Multi-Generator Ensemble Framework for Text-to-SQL.
+arXiv preprint arXiv:2411.08599 (2024).
 - Giray (2023) Louie Giray. 2023. Prompt engineering with ChatGPT: a guide for academic writers. Annals of biomedical engineering 51, 12 (2023), 2629–2633.
@@
 670–681.
+- Hui et al.
+(2024) Binyuan Hui, Jian Yang, Zeyu Cui, Jiaxi Yang, Dayiheng Liu, Lei Zhang, Tianyu Liu, Jiajun Zhang, Bowen Yu, Kai Dang, et al.
+2024.
+Qwen2.
+5-Coder Technical Report.
+arXiv preprint arXiv:2409.12186 (2024).
+- Hundman et al.
+(2018) Kyle Hundman, Valentino Constantinou, Christopher Laporte, Ian Colwell, and Tom Soderstrom.
+2018.
+Detecting spacecraft anomalies using lstms and nonparametric dynamic thresholding.
+In Proceedings of the 24th ACM SIGKDD international conference on knowledge discovery & data mining.
 - Ittoo et al. (2016) Ashwin Ittoo, Antal van den Bosch, et al. 2016. Text analytics in industry: Challenges, desiderata and trends. Computers in Industry 78 (2016), 96–107.
@@
 In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing.
-6769–6781.
 - Kaur et al.
@@
 International Journal of Transcontinental Discoveries 6, 1 (2019), 29–34.
+- Ke et al.
+(2024) Wenjun Ke, Ziyu Shang, Zhizhao Luo, Peng Wang, Yikai Guo, Qi Liu, and Yuxuan Chen.
+2024.
+Unveiling factuality and injecting knowledge for LLMs via reinforcement learning and data proportion.
+Science China Information Sciences (2024).
 - Kim et al.
@@
 In Forty-first International Conference on Machine Learning.
-- Klisarova-Belcheva et al. (2017) S Klisarova-Belcheva, G Ilieva, and T Yankova. 2017. BUSINESS INTELLIGENCE AND ANALYTICS–CONTEMPORARY SYSTEM MODEL. Trakia Journal of Sciences 15, 1 (2017), 298–304.
+- Kurt Schlegel (2023) David Pidsley Julian Sun Georgia O’Callaghan Christopher Long Kevin Quinn Fay Fei Edgar Macari Jamie O’Brien Kurt Schlegel, Anirudh Ganeshan.
+2023.
+Magic Quadrant for Analytics and Business Intelligence Platforms.
+https://www.gartner.com/doc/reprints?id=1-2HW1JC8Q&ct=240620
 - Lee et al. (2022) Chee Sun Lee, Peck Yeng Sharon Cheang, and Massoud Moslehpour. 2022. Predictive analytics in business analytics: decision tree. Advances in Decision Sciences 26, 1 (2022), 1–29.
+[…]
+- Li et al. (2024b) Chaofan Li, Yingxia Shao, and Zheng Liu. 2024b. SEA-SQL: Semantic-Enhanced Text-to-SQL with Adaptive Refinement. arXiv preprint arXiv:2408.04919 (2024).
 - Li and Jagadish (2014) Fei Li and H.
@@
 - Li et al.
-(2024) Haoyang Li, Jing Zhang, Hanbing Liu, Ju Fan, Xiaokang Zhang, Jun Zhu, Renjie Wei, Hongyan Pan, Cuiping Li, and Hong Chen.
-2024.
+(2024c) Haoyang Li, Jing Zhang, Hanbing Liu, Ju Fan, Xiaokang Zhang, Jun Zhu, Renjie Wei, Hongyan Pan, Cuiping Li, and Hong Chen.
+2024c.
 CodeS: Towards Building Open-source Language Models for Text-to-SQL.
@@
 In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023.
+- Li (2024) Jian-Zhong Li. 2024. opengauss: An open-source database for the era of artificial intelligence. Journal of Computer Science and Technology (2024).
+- Li et al.
+(2019) Zeyan Li, Chengyang Luo, Yiwei Zhao, Yongqian Sun, Kaixin Sui, Xiping Wang, Dapeng Liu, Xing Jin, Qi Wang, and Dan Pei.
+2019.
+Generic and robust localization of multi-dimensional root causes.
+In 2019 IEEE 30th International Symposium on Software Reliability Engineering (ISSRE).
 - Lian et al.
@@
 Journal of Computer Science and Technology Studies 6, 1 (2024), 225–232.
+- Liu et al.
+(2024) Xinyu Liu, Shuyu Shen, Boyan Li, Peixian Ma, Runzhi Jiang, Yuyu Luo, Yuxin Zhang, Ju Fan, Guoliang Li, and Nan Tang.
+2024.
+A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?
+arXiv preprint arXiv:2408.05109 (2024).
+- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A unified approach to interpreting model predictions. Advances in neural information processing systems 30 (2017).
 - Luo et al.
@@
 In NeurIPS 2024 Third Table Representation Learning Workshop.
+- Meduri et al.
+(2021) Venkata Vamsikrishna Meduri, Abdul Quamar, Chuan Lei, Vasilis Efthymiou, and Fatma Ozcan.
+2021.
+BI-REC: guided data analysis for conversational business intelligence.
+arXiv preprint arXiv:2105.00467 (2021).
 - Narang and Chowdhery (2022) Sharan Narang and Aakanksha Chowdhery. 2022. Pathways language model (palm): Scaling to 540 billion parameters for breakthrough performance. Google AI Blog (2022).
@@
 Designing business intelligence (BI) for production, distribution and customer services: a case study of a UAE-based organization.
-Business Process Management Journal 27, 4 (2021), 1275–1295.
+Business Process Management Journal (2021).
 - Popescu et al.
@@
 - Pourreza and Rafiei (2024) Mohammadreza Pourreza and Davood Rafiei. 2024. DTS-SQL: Decomposed Text-to-SQL with Small Large Language Models. arXiv preprint arXiv:2402.01117 (2024).
+- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog (2019).
 - Rane et al.
@@
 Techniques, and Opportunities (2024).
+- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych.
+2019.
+Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks.
+In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP).
 - Rivera (2024) Kevin Mejía Rivera.
@@
 1–4.
-- Robertson and Sparck Jones (1976) SE Robertson and K Sparck Jones. 1976. Relevance Weighting of Search Terms. Journal of the American Society for Information Science 27, 3 (1976), 129–46.
+- Robertson and Sparck Jones (1976) SE Robertson and K Sparck Jones. 1976. Relevance Weighting of Search Terms. Journal of the American Society for Information Science (1976).
 - Sun et al.
@@
 arXiv preprint arXiv:2306.00739 (2023).
+- Sun et al.
+(2019) Yu Sun, Shuohuan Wang, Yukun Li, Shikun Feng, Xuyi Chen, Han Zhang, Xin Tian, Danxiang Zhu, Hao Tian, and Hua Wu.
+2019.
+Ernie: Enhanced representation through knowledge integration.
+arXiv preprint arXiv:1904.09223 (2019).
 - Talaei et al.
@@
 63–68.
-- Team (2024) Qwen Team. 2024. Qwen2.5: A Party of Foundation Models. https://qwenlm.github.io/blog/qwen2.5/
-- team ([n.d.]) StarRocks team. [n.d.]. StarRocks. https://github.com/StarRocks/starrocks
+- Team (2024) Claude Team. 2024. Claude3. https://anthropic.com/claude
+- Team et al.
+(2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al.
+2023.
+Gemini: a family of highly capable multimodal models.
+arXiv preprint arXiv:2312.11805 (2023).
+- team (2022) OpenAI team. 2022. Introduction of OpenAI Text Generation APIs. https://platform.openai.com/docs/guides/text-generation
+- team (2023a) OpenAI team. 2023a. OpenAI CodeX. https://openai.com/index/openai-codex
+- team (2023b) StarRocks team. 2023b. StarRocks. https://github.com/StarRocks/starrocks
 - Wang et al.
-(2023) Bing Wang, Changyu Ren, Jian Yang, Xinnian Liang, Jiaqi Bai, Qian-Wen Zhang, Zhao Yan, and Zhoujun Li.
-2023.
+(2023b) Bing Wang, Changyu Ren, Jian Yang, Xinnian Liang, Jiaqi Bai, Qian-Wen Zhang, Zhao Yan, and Zhoujun Li.
+2023b.
 MAC-SQL: A Multi-Agent Collaborative Framework for Text-to-SQL.
@@
 - Wang et al.
-(2020) Bailin Wang, Richard Shin, Xiaodong Liu, Oleksandr Polozov, and Matthew Richardson.
-2020.
+(2020a) Bailin Wang, Richard Shin, Xiaodong Liu, Oleksandr Polozov, and Matthew Richardson.
+2020a.
 RAT-SQL: Relation-Aware Schema Encoding and Linking for Text-to-SQL Parsers.
@@
 arXiv preprint arXiv:2402.05672 (2024).
+- Wang et al.
+(2025) Xin Wang, Zirui Chen, Haofen Wang, Leong Hou U, Zhao Li, and Wenbin Guo.
+2025.
+Large language model enhanced knowledge representation learning: A survey.
+Data Science and Engineering (2025).
+- Wang et al.
+(2023a) Yizhong Wang, Yeganeh Kordi, Swaroop Mishra, Alisa Liu, Noah A Smith, Daniel Khashabi, and Hannaneh Hajishirzi.
+2023a.
+Self-Instruct: Aligning Language Models with Self-Generated Instructions.
+In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
+- Wang et al. (2020b) Yaqing Wang, Quanming Yao, James T Kwok, and Lionel M Ni. 2020b. Generalizing from a few examples: A survey on few-shot learning. ACM computing surveys (csur) (2020).
 - Wei et al.
@@
 24824–24837.
+- Weir et al.
+(2020) Nathaniel Weir, Prasetya Utama, Alex Galakatos, Andrew Crotty, Amir Ilkhechi, Shekar Ramaswamy, Rohin Bhushan, Nadja Geisler, Benjamin Hättasch, Steffen Eger, et al.
+2020.
+Dbpal: A fully pluggable nl2sql training pipeline.
+In Proceedings of the 2020 ACM SIGMOD International Conference on Management of Data.
 - Wieder and Ossimitz (2015) Bernhard Wieder and Maria-Luise Ossimitz.
@@
 Procedia Computer Science 64 (2015), 1163–1171.
-- Yu et al.
-(2018) Tao Yu, Rui Zhang, Kai Yang, Michihiro Yasunaga, Dongxu Wang, Zifan Li, James Ma, Irene Li, Qingning Yao, Shanelle Roman, Zilin Zhang, and Dragomir R.
-Radev.
-2018.
-Spider: A Large-Scale Human-Labeled Dataset for Complex and Cross-Domain Semantic Parsing and Text-to-SQL Task.
-In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing.
-3911–3921.
+- Wu et al.
+(2021) Kun Wu, Lijie Wang, Zhenghua Li, Ao Zhang, Xinyan Xiao, Hua Wu, Min Zhang, and Haifeng Wang.
+2021.
+Data augmentation with hierarchical SQL-to-question generation for cross-domain text-to-SQL parsing.
+arXiv preprint arXiv:2103.02227 (2021).
+- Yang et al.
+(2024) An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, et al.
+2024.
+Qwen2.
+5 technical report.
+arXiv preprint arXiv:2412.15115 (2024).
+- Yao et al.
+(2023) Shunyu Yao, Jeffrey Zhao, Dian Yu, Nan Du, Izhak Shafran, Karthik Narasimhan, and Yuan Cao.
+2023.
+React: Synergizing reasoning and acting in language models.
+In International Conference on Learning Representations (ICLR).
 - Zeng et al.
@@
 161–172.
+- Zhang et al.
+(2023b) Baoli Zhang, Haining Xie, Pengfan Du, Junhao Chen, Pengfei Cao, Yubo Chen, Shengping Liu, Kang Liu, and Jun Zhao.
+2023b.
+ZhuJiu: A Multi-dimensional, Multi-faceted Chinese Benchmark for Large Language Models.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations.
+- Zhang et al.
+(2024) Jiayu Zhang, Jiangrong Shen, Zeke Wang, Qinghai Guo, Rui Yan, Gang Pan, and Huajin Tang.
+2024.
+SpikingMiniLM: energy-efficient spiking transformer for natural language understanding.
+Science China Information Sciences (2024).
+- Zhang et al.
+(2023a) Yi Zhang, Jan Deriu, George Katsogiannis-Meimarakis, Catherine Kosten, Georgia Koutrika, and Kurt Stockinger.
+2023a.
+Sciencebenchmark: A complex real-world benchmark for evaluating natural language to sql systems.
+arXiv preprint arXiv:2306.04743 (2023).
+- Zhou et al. (2024) Xuanhe Zhou, Zhaoyan Sun, and Guoliang Li. 2024. Db-gpt: Large language model meets database. Data Science and Engineering (2024).